A Bibliography of Publications in ACM SIGNUM

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

18 November 2023 Version 1.49

Title word cross-reference

(10^{31} − 1)/9 [Hig86]. (k, p) [DJ79]. 1 [Rol90]. 2 [PP75, Rol90]. = [Kre95, LK96]. A [WB79]. δ^{2} [Jon82]. LU [Zah89]. n [Mor69, Str77]. N + 1 [DZ88b]. N × N [Par69].

-ary [Mor69]. -stable [WB79]. -th [Str77].

1 [DPS92, Hop97, Ile85, Tem83]. 1/96 [Dry96]. 16-18 [Ano73q]. 1971 [Fri72a, Kah72]. 1973 [Ano73b, Ano73f, Ano73i, Ano73a, Ano73j, Ano73q, Ano73n, Ano73m, Ng73]. 1974 [Ano73g, Ano73d, Ano73r]. 1983 [Smi83b]. 1995 [BD95].

2 [Ano73k]. 2-by-2 [Mol74]. 205 [Tem83].

3 [Smi83a]. 360 [CCHT67]. 360/370 [Boy75, TC75].

6600 [DM73]. 6600/7600 [DM73].

7000 [Nik72]. 74 [Ano73h]. 754 [Tyd94]. 7600 [DM73]. 77 [Rei80].

80th [Gil98]. 8th [Ano73a]. 8x [Smi83b, Smi88a, Smi88b].

90 [DS94, IFI93]. ’95 [Ano95f].

A-stability [Wan79]. absolute [Wis78]. Abstract [Boy79, Ber75, CW75a, EHHM75, ES75a, ES75b, Joh75, Jon75, LS75, Mad75, NDM75, PP75, Pri75, Rup75, ST75, SW75, SS75]. Abstraction [Smi98]. Abstracts [Ano93a]. academic [Kel76]. Accelerating [McN98]. Accuracy [Kre96]. accurate [Dun90c].

Achieving [Hod95]. ACM [Ano73b, Ano73k, Ano93b, Han77, RS97]. ACM-SIGNUM [Han77]. Activities [Jon79, Ein78]. activity [Voi81]. Ada [BB91]. adaptation [Str76]. adaptive [Ric73, dD78]. Addition [RBSF93]. additions [Han70]. Advanced [Jon79]. AFWL [MND75, NDM75]. agreements [BFF83]. Aided [FK96, Fra90]. Aitken [Jon82]. Algebra [Ano95e, DL85a, DL85b, DCHH85, Ste96a, Ste96b, DCDH87, Far74, Far76, HKL73, LE85, WG80]. Algebraic [Kre95, LK96, Law74]. Algorithm [DK72, See80, GMSW85, Ric73, Rol87, Str77, Tan75, Tee80, UF69, Zah89, dD78]. Algorithmic [Ric76b]. Algorithms [Ano73q, Bra95, For73, Hor74, Dun77, Gup72, KCD90, dP87]. Almost [Jer89]. alternative [GP87, Kni87, WB79]. amendments [FRS77]. AMS [Ano73c, Ng73]. Analysis [Ano93c, Car92, Dun89b, FKK96, Kah91, McK81, Ano71, Cha87, Han68, Mol68b, Mol68c, Mol69b, Mol69c, Ort66, RW71a, Ric79, Smi88b, Ste80, Tra66, Tra67a, Tra67b, Tra67c, Vol79]. Analytic [Rol97]. animation [Cyn88]. annotated [EJ76]. Announcement [Ano95f, Ano95e, Ano95a, Smi83a]. Annual [Ede91]. ANS [FRS77, Kni76a, Kni76b]. Any [Jer89]. anyone [Par67]. aphorisms [Kah68]. Application [KS97, Pri75]. Applications [Ano73m, Dun91, FKK96, Gup72]. Applied [BD95, DW95, DW96, Joh75]. approach [Bra79, KM83]. approaches [Smi79]. approximate [MC97]. Approximately [Kre96]. approximating [Law68]. Approximation [DZ88a, Dun91, Dun76b, Tra67b]. approximations [Dun87a, Dun87b, Dun88]. Apr. [Kah72]. April [Ano73k, Ng73, Rob73]. arbitrary [Hua78]. Architecture [Ris76, LS75]. area [Ric81a]. areas [Ric76a]. Argonne [Ash69].

Arguments [Wan78]. Arithmetic

[BB91, CCG^{+}85, CKP^{+}79, Fel79c, Rei79b, Cle84, Cor83, CY79, Dit93, Dun89a, Hul79, Moo80, PSW90, IFI92, Rei93]. arrays [EL81]. art [Mol69d, Tra68]. ary [Mor69]. association [Rei80]. Astrogeometry [FKK96]. Atlanta [Ano73b]. AUGMENT [CY79]. August [Ano73b, Ano73a, Smi83b]. Austin [Ano73i, Ano73k]. automated [Ard80]. Automatic [BD95, Jer89, Pfe87, SB82, ES75b, dD78]. Availability [FC84, Smi83a]. Available [Ano91, Don92, Cod88, DM76]. Averaging [WK96]. AVIRG [Kob92].

Background [Dun76a, Dri82]. Band

[EL81]. banded [Bra90]. Basic

[DCHH85, BF79, DCDH87, Law74, Kah81b, DL85a, DL85b]. basics [Cod80]. bc [Dit93]. be [Ard80, Dun90c]. Beale [Ano86]. believe [Hul73]. Best [Dun91]. better [Ske79a, WK96]. Bibliography [Uso70, EJ76]. big [Gro91]. Binary [CKP^{+}79, PS79, Rol97]. Biofeedback [Kos96]. Birthday [Gil98]. BLAS [Aha85, DDHH86, DDCH89, Hop97]. Boeing [PP82]. Bolts [Smi98]. Book [Ano74a]. Books [Ano98c, Ano98g]. bordered [GP87, Kni87]. boundary [DM76, SAW85]. boundary-value [DM76]. Bounded [WK96]. brief [Fri72a, GN81].

C [Bra90]. cache [Str81]. calculation [Bro73, Tra68]. calculations [Bus87, EHHM75, Nik72, Ste81]. calculators [Ste81]. Calendar [Ano73e]. Call [Ano73h, Ano93b, Ano93c]. Can [Kre96, Ard80]. Cancellation [Del96, Duk96, Dry96]. capability [CS83]. Carnegie [Ano73q]. Carnegie-Mellon [Ano73q]. Case [FK96, Mil79]. Cases [Dun89b]. caused [Cyn90]. causes [Tyd94]. CDC [DM73, Nik72]. Center [ADHR73, Ano73m, KM83, But79b]. central [Cyn87]. Certification [Uso70, Ash69, Uso69]. Certifying [Han69]. chain [Law88]. Chair [Ano98e]. challenge [Dun86a, Kar85]. CHAMP [EHHM75]. Chan [GP87]. change [SAW85]. chapter [Gla85]. Characteristics [Cod71]. Chebyshev [Dun76b, Mac93]. Choosing [FK96]. City [Ano73n]. Class [AA96, But79a, Str76]. classification [BHK83, Bol75]. Code [DM76, EHHM75, ES75b, MND75]. code-codes [MND75]. codes [Joh75, MND75, Say80, Ske79a, Ste79]. coefficients [Bro73, Cyn87]. Combatting [Lin81]. commands [Dit93]. Comment [HS93]. Comments [Fel79a, IFI92]. Committee [Jon79, Uso69]. commonly [Ker72]. communications [Ano74b]. Commutative [RBSF93]. Comparative [Wit67]. Comparison [McN98, CPR69, EH79, Eva89, KCD90, Smi79]. Comparisons [Rol97, Rol90, Tem83]. Compatible [IFI92, PSW90]. Compilation [BD95]. Complex [Rei80, Dun89a, Wis78]. Complexity [Ano73c, Ano73q, Hor74]. Complicated [Kos96]. composite [PP75]. compositions [Mor69]. comprehensive [Mol68a]. Computation [Cod71, GOP^{+}79, GOP^{+}80, BF79, Cyn87, Dun72, GN88, Gre79, Hul79, Kul83, KM83, Ric76a, RGO^{+}79, Ric81b, Smi83c, Xan86]. Computational [Ano73c, Ano73f, Ano73i, Cod80, Zah89]. Computationally [Kos96]. Computations [Bai98, Kos96, KK96, KS97, NK97, Bat81, Mol69d, Ste68]. Computer [ADHR73, Ano71, Ano73d, Ano73j, BB91, Car92, FK96, Kos96, LK96, Rei79b, Ste96a, Ard80, Bol75, Cha87, Fra90, Law68, PP82, Wit67]. Computer-Aided [FK96, Fra90]. computers [Bat81, Don84, DM73, LS75, Nas81, Nik72, Str81]. Computing [Ano73k, DW96, Gil98, Hop97, Jon79, CG89, Dun76a, GAA^{+}86, Wis78, DW95]. concept [Joh75]. concerning [Kah68, Kni76b, Sor81]. conditioned [Tee72]. conditioning [Hua76]. Conference [ACM79, Ano73b, Ano73e, Ano73f, Ano73g, Ano73d, Ano73i, Ano73j, Fos91, Pow73, Ano79, Ano73l, Ano93b, Ano95a, Ano96f]. Congress [Ano73h]. conjecture [Far74, Far76]. connected [Kro72]. Consortium [McK81]. constants [ABB^{+}74]. Constrained [Ano73r]. constraints [Kro87]. Construct [IFI93]. construction [EH77]. continuous [Dun76b, Law68]. continuum [SW75]. Contribution [Kah81a]. Control [NK97, But73, Ehl74, Fra81]. Convention [BD95]. Convergence [McN98]. conversion [PV83a, PV83b]. core [Bra79]. corner [Mol69a]. correctness [Dun86b]. corrector [Bro73]. Correspondence [Eng74]. cos [Tur69]. COSERS [Ard80]. COSMIC [Ano91]. counting [Mor69]. course [Cli82]. Cramer [Dun80, Mol74]. Cray [Tem83]. criteria [But73, Ehl74]. Cubpack [MC97]. Current [Ano98a, Ano98b, SAW85]. curves [Law68]. Cyber [Tem83].

D [PP75]. data [Dun90b, Ein86, Law68, Pay86, Ric76a, Sch79a]. Dates [Bra95]. dc [Dit93]. DCG [ES75b]. Dec [Kah72]. Decimal [Ris76, PV83a, PV83b]. Deck [Hod95]. Decomposing [Pro88]. decomposition [Dou92, Gol78, Ham85, Zah89]. definite [Bra90]. definition [Fri72b]. deflation [GS74, GP87]. Degree [PH83a, DJ79]. Delaney [Dry96]. Denmark [DW96]. Denormalized [FW79]. Dense [Ede91]. Department [Ano71, FJ79]. Departments [Ano98b, Ano98c]. dependent [ABB^{+}74]. Derivation [Rol97]. Derivatives [AA96]. description [GN81]. Design [Chi85, FK96, Fox79, Boy75, Fra90, Kah68, Nas90, TK97, TC75]. Desirable [Cod71, Hul79, Dun80, GN88]. Determination [Whi94]. determining [Lin81]. Detroit [Ano73d]. Development [ACM79, Ano73p, AK79, Ano79, CY79, Kel76, NDM75, dP87]. developments [Rup75]. differ [Pro88, Zah89]. difference [Cyn87, Cyn88]. Differential [Ano75, Ske79b, BL73, But79a, Chi86, Coo79, Cur73, Dah79, Fuj79, Gla85, GS75, HB76, Hin80, Kro72, Mad75, Ric76b, Str76, Str77, SS75, War76]. Differentiation [BD95, FK96, Jer89, Pfe87]. Difficulties [Dun90b]. Difficulty [Tur69]. DIFSUB [SG72]. dimensional [OR75, SW75, WK95]. direct [Xan86]. directions [RGO^{+}79]. DIS [Rei93]. discrete [Law68]. disguised [Ful86]. displacement [Pri75]. disquieting [Kni77]. distributed [Gar86]. Distribution [DG85, DK72, Ile85]. documentation [Ost79]. DOE [Jon79]. domain [Dou92]. Double [Fig95, Dun89a, MC97, Rei80]. DPUP [Gar86]. Draft [CKP^{+}79, Kni76a, PS79, FRS77, IFI92, Kni76b, Rei93]. drawn [Dri79]. dynamic [CS83]. dynamics [MND75, NDM75].

easy [MC97]. Editor [Law73, Kni76b]. Editorial [Ano98d, Ano98f]. education [TK97]. Educational [RS97, Ric73]. effects [Lin81]. efficiency [HKL73]. Efficient [ES75a, SS75, ST75, Tee80]. effort [Ric79]. EFL [Fel79b]. eigenproblems [Str81]. eigensystem [Gol78]. Eigenvalue [RW71b, Nik72]. Eigenvalues [Par67, EH77]. Eigenvector [RW71b, Fra81]. eigenvectors [Tee74]. EISPACK [Smi83a, HS93]. Election [Ano73o]. electronic [DG85, Ano93c]. elegant [EL81]. element [CW75a, Pri75]. elementary [Hul79, Tho87]. elements [Dun89a, Rol88, Zah89]. Eliminate [FW79]. elimination [CG85, Cyn87, DEGR88, ES75a, Eva89, Mac89, Tre85]. elliptic [SS75]. ELLPACK [Ric79]. empty [dB90a, dB90b]. Enable [IFI93]. ENEP [SW75]. Energy [But79b, FJ79, KS97]. engineering [Ard80, Smi83c]. England [Ano73f, Ano73g, Ano73r]. enhance [Dri82]. Enhanced [Cor83]. enquiries [RW71a]. Environment [ACM79, BF79, Fel79a, Kah81a, RW71a, Sch79b, AK79, Ano79, Don84, For78, GAA^{+}86, Nas90, Sch79c, Wil86]. Environments [Fos91, Ric96]. Epsilon [HS93]. equal [EH77]. equality [Dun86b]. equation [Chi86, Cur73, Cyn87, Han69, Hin80, War76]. Equations [Ske79b, BL73, But79a, Coo79, Cyn88, Dah79, Don84, DZ88b, Ehl74, GN81, Gla85, GS75, HB76, Kro72, Law72, Mad75, Rhe69, Ric76b, Str76, Str77, SS75, Tan75, Ano75]. equilibration [Dun80]. Errata [Far75, Rob73]. Error [Del96, Dit93, Duk96, FKK96, Fra81, Gre79, CG85, Coo79, Cyn90, DM73, Dry96, Han68, Kni77]. Error-free [Gre79]. Errors [WK96]. Essex [Ano73f]. Estimates [WK96, Coo79]. estimating [CG85]. Estimation [FKK96, Kle78, Sor81, Tra67b]. ETNA [Ano93c]. evaluating [UF69]. Evaluation [Ash69, Ric91, Wan78, Dun86a, Dun90a, Kro71, Kro72]. Evanston [Ng73]. exact [DZ88a, Zie81]. exam [Ano71, For70]. Examination [Tra67a, Mol68a, Mol68b, Mol68c, Mol69b, Mol69c, Tra66, Tra67c]. Example [Del96, Duk96, Dry96, Han68, Sor81, Tee72]. Exception [IFI93]. exceptions [ELP79]. Excerpts [Mol68a]. exchange [Sny79]. exercise [dB90a, dB90b]. existence [DJ79]. expansion [DK72]. Expansions [Mac93]. experience [Du 79]. Experiences [Dri79, OC70]. experiments [Lew83, SG72, Wis78]. explicit [Ker72]. Exploration [NK97]. explosions [EHHM75]. exponential [Str76]. Extended [Boy79, DDHH86, DCHH85]. Extending [Boy79]. extension [DL85a, Gla85, Kul83]. extensions [DL85b, Dri82]. extrapolation [dD78].

F [Hor74, DK72]. facilities [Ric81b]. factor [Hig86]. Factored [WB79]. factorization [Ker72]. Factors [Kel76]. Fall [Ano73n]. family [Joh75]. Fast [GLJ^{+}91, Nik72, Han73, Rol87, Chi85]. Feasibility [Dun90a]. February [Ano73d]. File [Ste91]. Files [Gro92, Gro91]. Finding [Bai98, Che97]. Finite [CW75a, Cyn90, Gre79, Pri75]. First [AA96, Ede91, Str76]. fitting [Dun90b, Law68]. Fixed [Bri94]. Floating [Cod79, CCG^{+}85, CKP^{+}79, Fel79c, PS79, Rei79a, Ris76, RBSF93, Tyd94, BF79, Cle84, ELP79, Hul79, Kah68, KP79, Pay86]. Floating-Point [Rei79a, Ris76, RBSF93, Tyd94, CCG^{+}85, BF79, Hul79, KP79]. fluid [MND75, NDM75]. form [Par69]. formatting [Dri79]. formula [ES87]. Forsythe [Mol72]. FORTRAN [FRS77, Boy75, CW75b, Kni76a, Kni76b, Kul83, Ske79a, Smi83b, SS75, TC75, Bra79, CCHT67, Cor83, Don84, DCHH85, DS94, EL81, Eng74, Hul73, IFI93, Pre79, Rei80, Smi88a, Smi88b, Wag84]. Fortran-90 [DS94]. forward [Cyn88]. Foundations [TK97]. Four [Kah68]. fractions [PV83a, PV83b]. fracture [Xan86]. free [Gre79]. fronts [Pri75]. FTP [Ste91]. Full [Hod95]. Function [Nas81, Dun86a, Dun90a, Tho87]. Functions [Bur73, Mac93, Ng73, PH83a, PH83b, Rei79a, Wan78, BF79, Dun76b, Hul79, Law68, WK95]. Further [Gol78]. Future [GOP^{+}79, GOP^{+}80].

Games [Kos96]. GAMS [BHK83, BKM88]. Gaussian [CG85, Cyn87, DEGR88, ES75a, Mac89, Tre85]. Gear [DJ79]. general [But79a]. Generalization [Zie74]. Generalizations [Zie81]. Generalized [Ano73m, Zie78a, Zie78b]. generating [See80]. generation [CW75a, ES75b, Jon75]. Generations [Rol88]. generator [Ful87]. geometric [Fra90]. George [Mol72]. German [Gil98]. Givens [GLJ^{+}91, Han73, Hop97]. Goes [GLJ^{+}91]. graphic [SW75]. Group [Ein73, Fel79a, Han77, For73, Jon79]. guide [BKM88]. Guidelines [UF69]. Gulf [Kos96].

H [CCHT67]. Hamburg [BD95]. hand [Ste81]. hand-held [Ste81]. Handling [ELP79, IFI93]. hard [Kro87]. Hardware [Cod71, Dun72, Kah68]. held [Ste81]. Help [Car74]. Hermitian [Gol78, Str81]. Hessian [Sor81]. High [Fos91, Ris76]. High-Level [Fos91, Ris76]. Hilton [Ano73d]. Hirondo [Kah72]. Horner [Dun89b]. hosts [DW96]. Hotelling [GS74]. Householder [Rol90]. HULL [MND75]. hydrocode [NDM75]. hyperbolic [ES75b]. hypertext [Nas90].

IBM [Boy75, CCHT67, KM83, TC75]. ICASE [Voi81]. Idea [Kos96]. Identities [Bai98]. IEC [IFI92]. IEEE [Cod79, Tyd94]. IEEE-754 [Tyd94]. If [Kre96]. IFIP [Ano73h, Ano95a, Ein73, Fel79a, Fos91]. II [Dun87b]. III [Dun88]. ill [Tee72]. ill-conditioned [Tee72]. Illinois [Add81]. illustrative [Han68]. immiscible [Pri75]. Impact [Cod79, Fel79c]. Implementation [Chi97, Chi85, ES75a]. implementations [Cle84]. Implementing [Tho87, GMSW85]. implicit [Eva80]. important [Xan86]. Improvement [Dun89a]. Improving [HKL73]. IMSL [AK79]. incomplete [Ske79a]. Independent [Rei93, CCG^{+}85]. Industrial [BD95, McK81]. inequalities [Tee80]. Information [Ano73o]. Initial [OC70, ES87, Hin80]. Innocuous [Fig95]. Input [Wan78]. Instability [Mac93]. Institutional [Ano95b, Ano95c, Ano95d, Ano96a, Ano96b, Ano96c, Ano96d, Ano98a, Ano98b].

Instructions [Fel73]. integer [Rol87]. integrals [MC97]. integrating [Kro72]. Integration [AA96, CPR69, WK95, dD78]. integrators [BBG88]. Intelligent [NK97]. interactive [MC97, SW75]. Interesting [Hod96, Lyn68]. interface [Boy75, MC97, TC75]. interfaces [Rup75]. International [Ano73i, Ano73a, BD95, Rei93, Ano95e]. Interpolation [Aki96, DZ88a, FC84]. Interval [Cle84, Kos96, KK96, KS97, Kre95, LK96, NK97, WK96, Moo80]. inverse [Ste75]. Inverses [Ano73m, Zie78a, Zie78b, Zie81]. inversion [Eva80, Fra81]. inverting [GS74, Zie73]. Iowa [Ano73n]. Isn’t [RBSF93]. ISO [IFI92]. ISO/IEC [IFI92]. ISSAC [Ano95f]. Issues [BFF83, DL85a, GMSW85]. Iterative [Che97, Zie78a, DT76]. IV [CCHT67]. IVPS [Ske79a].

J [Dry96, Hor74, KM83]. Jacobian [OR75]. January [Ano73g, Ano73r]. Jensen [Kni72]. Julian [Bra95]. July [Ano73f]. June [Ano73j]. justification [Dri82].

KFA [Hak85]. Khachian [Tee80]. Known [Kre96]. Konrad [Gil98]. Kuki [Kah72]. Kutta [ES87].

Laboratories [CM78]. Laboratory [Ano73r, Fri73]. Lagrange [DZ88a]. Lagrange-type [DZ88a]. Language [IFI92, Jer89, Jon79, Aha85, Dri82, Ein86, Fel79b, PSW90, Ric79, Ric81b, Wag84, Rei93]. Languages [Fel79c, Ris76, CG79, Ric76a, Smi83c]. LAPACK [Don92]. Large [Ede91, Wan78, Boy75, Kel76, MND75, Say80, Str81, TC75]. Lawrence [Fri73]. LCAS [Kah91]. least [Bra87, KCD90, Ste75]. leg [Dah79]. legitimate [Ric81a]. length [CCG^{+}85]. Letter [Ano98e, Kni76b]. Letters [Ano98f]. Level [Fos91, Hop97, Ris76, DCDH87, DDCH89]. LIA [DPS92, Rei93]. LIA-1 [DPS92]. Libraries [Ein73, Bol75, BFF83]. Library [ADHR73, Ano91, FJ79, RW71b, BBG88, CCHT67, Fox78, Fox79, Gla85, Tho87, VH82, Du 79, FB77]. license [BFF83]. limiting [Kel76]. Linear [Ano95e, DL85a, DL85b, DCHH85, Ede91, Coo79, Dah79, Don84, DCDH87, Far74, Far76, GN81, Han69, HKL73, Law72, Law74, LE85, Ste75, Tan75, Tee80, WG80, WB79, Whi94]. lines [Mad75, WB79, Xan86]. list [Ske79a]. Livermore [Fri73]. Local [Aki96, Coo79, Rad73, Han77]. Lost [Car74]. Loughborough [Ano73l, Pow73]. LSODE [Hin80]. LSODI [Hin80].

MACHAR [Cod88]. Machine [HS93, ABB^{+}74, FB77, Smi79]. MACSYMA [CM78]. Madison [Ano73m]. mail [DG85]. Making [Kro87]. management [Sny79]. Manipulating [Rei79a]. Martin [Ano86]. Mass [Kre96]. Masses [Kre96]. material [Rup75]. MATH77 [Ano91]. Mathematica [Ano96f]. Mathematical [ADHR73, Ano73a, Ano73l, Ano73p, Bai98, Bat75, Hak85, Pow73, BHK83, Boy79, Chi86, Cli82, DG85, Dri82, EJ76, Kel76, Kro71, Nas90, Ric81a, Smi79, VH82, Wil86, Wit67]. mathematician [Fri72b]. Mathematics [Ano73g, Ano73m, BD95, Cod80, Fri72a, Fri73, PP82, TK97]. MATLAB [GLJ^{+}91]. matrices [EH77, Eva89, GS74, Hua76, Hua78, Ker72, Par69, Pro88, Tee74, Zah89, Zie73, Zie78b]. Matrix [DS94, Ste81, DGLP82, Eva80, Fra81, GN84, GN88, Gol78, Jen71, Mol69d, See80, Tee72, Zie73, Zie74, Zie78a, Zie81]. Matrix-Matrix [DS94]. matters [Kro72]. maximum [DJ79]. May [Ano73k, Ano73q]. McMaster [RW71b]. Measure [Kre96]. mechanical [SW75]. Mechanics [Ano73i, Xan86]. Meeting [Ano73n, Ano75, Ske79b]. Mellon [Ano73q]. memorial [Ano86]. memoriam [Kah72]. memory [GN88, Mol72, Str81]. mesh [CW75a, CS83, Jon75]. Method [Car74, Dun89b, FK96, McN98, Mad75, ST75, WB79, Xan86]. Methods [Aki96, Ano73i, Ano73r, BL73, McN98, Ber75, But79a, Coo79, Dah79, DJ79, DT76, Eva89, Fuj79, Ful86, Kro71, Kro72, Lam79, Law68, OC70, Pri75, Say80, Smi83c, Str76, Wan79, WB79]. MGNet [Dou92]. micro [Bat81]. micro-computers [Bat81]. microcomputer [Ste80]. microcomputers [Chi85]. Microprogrammed [Moo80]. military [TK97]. minicomputers [Nas76]. minimization [Nas81]. Mission [Ano97]. Mixed [Ein86]. modelling [Xan86]. models [SAW85, Whi94]. Modern [BB91]. Modified [Car74, Hop97, GS74]. modules [Bra79]. monotone [Dun87a, Dun87b, Dun88, FC84]. Monotonicity [KK96]. Moore [Zie81]. motivations [Nas90]. Moving [SAW85]. Multi [Aha85, WK95]. multi-dimensional [WK95]. Multi-language [Aha85]. multidimensions [WB79]. multigrid [Dou92]. Multiple [McN98, RS97, Dit93]. multiple-precision-arithmetic [Dit93]. Multiplication [DS94]. multiplications [Par69]. multistep [Dah79, Str76, Wan79, WB79]. multivariate [Ham85]. mysteries [Tre85].

NAG [BBG88, Du 79, For73, FB77, Gla85]. Name [ABB^{+}74]. National [Ano73b, Ano73j, Ano73r, But79b]. nature [RGO^{+}79]. Navy [CM78]. Needed [NK97]. Netlib [Gro91, Gro92]. Network [FK96, Dou92]. News [Gro92, Gro91]. Newsletter [Dry96, Rob73]. Newton [Car74, McN98, OR75, Sor81]. NEXUS [GAA^{+}86]. Nice [Dun89b]. No [Law79]. non [Dun76b, Rol88, SB82, Whi94]. non-continuous [Dun76b]. non-linear [Whi94]. non-stiffness [SB82]. non-unique [Rol88]. nonacademic [Fri72a]. Nonconvergence [Che97]. Nonlinear [Ano73i, Rhe69, Ber75, DT76, DZ88b, ES87, KCD90]. nonsymmetric [ES75a]. Nordsieck [DJ79]. Normalized [PS79]. Note [Aki96, Boy75, Ful87, Hin76, Hua76, CG85, EH77, Jon82, KCD90, War76]. Notes [FJ79, Par69]. Notice [DDHH86, DDCH89]. Nottingham [For73]. number [Ful87, Gre79, Par69]. Numbers [FW79, Rei79a, PV83a, PV83b]. Numeric [Ste96b]. Numerical [ACM79, AA96, Ano73q, Ano73r, Ano93c, Bai98, Ber75, Car92, Cod79, Ein73, Fel79a, FK96, Ful86, GOP^{+}79, GOP^{+}80, Hor74, Kos96, Lew83, McK81, Nas76, PP82, RGO^{+}79, Ric91, Say80, Sch79a, Sch79b, Ske79b, Smi98, Ste96a, Tra67b, WG80, Ano71, Ano79, BFF83, CPR69, Cha87, CG79, Dun77, Dun90c, Ein78, Far74, Far76, For78, Fra81, Fri72a, Fri72b, Fri73, Fuj79, Gan76, Hul79, Kah81b, Kro71, Mad75, Mol68b, Mol68c, Mol69b, Mol69c, OC70, Ort66, RW71a, Ric76a, Ric81b, Rup75, Sch79c, SG72, ST75, Ste80, Str77, Tra66, Tra67a, Tra67c, TK97, Vol79, dP87, HAC^{+}85]. Nuts [Smi98]. Nuts-and-Bolts [Smi98].

objectives [Nas90]. observations [Ric76a]. Obstacles [Cha87]. occurring [Ker72]. October [Ano73n, Ano73m]. ODE [Hin76, Ste79, Xan86]. ODE-codes [Ste79]. ODES [Ske79a]. off [CS83]. off-the-shelf [CS83]. often [WK96]. OIL [Joh75]. OlympiODE [Add81]. one [Dah79, DZ88b, HB76, Pro88, ST75, SW75, Wan79, Zah89]. one- [Wan79]. one-dimensional [SW75]. one-leg [Dah79]. one-phase [ST75]. only [DZ88b, Pro88, Zah89]. operations [Law74]. Opinions [Kro72]. optimal [EH79]. Optimization [Ano73r, MTW80]. Order [AA96, Str76, Str77, War76]. Ordinary [Ske79b, But79a, Coo79, Cur73, Dah79, Gla85, HB76, Hin80, Kro72, Str77]. organizing [Han77]. oriented [Gup72]. orthogonal [Bra87]. Other [FKK96, Sny79]. output [Kni77]. overflow [Lin81].

package [Gar86, GN81, GN84, Jon75, PP75, Kle78]. packages [Boy75, CY79, Kel76, TK97, TC75, dP87]. Papers [Ano73h, Ano73k, Ano73l, Ano93b, Ano93c]. PARA95 [DW95]. PARA96 [DW96]. parabolic [PP75]. Parallel [Ano73q, DW95, DW96, Hor74, Dun86a, Eva89, GN88, Kar85, LS75, NDM75]. parameter [Whi94]. Parameterization [Fel79a, For78, Smi79]. parameters [BF79, Xan86]. part [Nas90]. Partial [Ano75, Mac89, Mad75, Ric76b, SS75]. participation [Kel76]. Particular [Kos96]. Pascal [Vol79]. Past [GOP^{+}79, GOP^{+}80]. patterns [Bat75]. Penrose [Zie81]. perfect [Dun90a]. Performance [CCHT67, Don84, Ric91]. Periodic [Wan78]. permutation [See80]. permutations [Rol88]. persymmetric [Gol78]. Perturbation [Dun89b, Ste75]. Ph.D. [Ano71, Ano93a, Ano96e, Mol68a, Mol68b, Mol69b, Tra66, Tra67c]. phase [ST75, SAW85]. phenomena [Cyn88]. philosophy [CG89]. Physical [Ano73r, Cyn88]. Pioneer [Gil98]. pipeline [LS75]. pivoting [CG85, ES75a, Mac89]. PL [Dri82]. PL/I [Dri82]. planar [CW75a]. plans [Gla85]. plea [Kro71]. Point [Bri94, Cod79, CKP^{+}79, Fel79c, PS79, Rei79a, Ris76, RBSF93, Tyd94, BF79, Cle84, CCG^{+}85, ELP79, Hul79, Kah68, KP79, Lam79, Pay86]. Poisson [Chi85]. Polynomial [Che97, McN98, DJ79, Tra68]. Polynomials [Ste96b, Lin81]. PORT [Fox78, Fox79]. Portability [But79b, DPS92, Dri82]. portable [Fox78, HKL73]. positive [Bra90]. possible [TK97]. Potential [KS97]. practicable [Zie78a]. practice [Du 79]. pre [Hua76]. pre-conditioning [Hua76]. precision [Dit93, Rei80, Sch79a]. precompiler [CY79]. preconditioning [EH79]. Preferences [Rei79b]. Present [GOP^{+}79, GOP^{+}80]. presented [Ano73k, Ano73l, BD95]. Principles [Rei79b, Fox79]. Probability [Bur73]. Problem [Fos91, Kah71, Mac93, Ric96, GAA^{+}86, Jen71, Kni72, Ste75]. Problems [Ano73f, Kos96, ML67, Bra87, Chi86, Cur73, DM76, DGLP82, ES75b, ES87, GS75, MND75, PP75, Pri75, ST75, SAW85, SW75, Wit67]. Proceedings [ACM79, Ano79, KM83, Ano75]. Process [Ano73c, Eva80]. processing [Gar86, Kar85]. Producing [Sch79b, Sch79c]. product [Par69]. Profile [Fri73]. Program [Ein73, Kah81a, RW71b, Ste91, Bol75, Bra90, BFF83, Dri79, MC97, OC70, PP75, SW75]. Programmable [Dri79]. programmers [Say80]. Programming [ACM79, Ano73a, Fos91, Kah81a, LE85, Ric79, Ano79, CW75b, Ein86, Fel79b, GMSW85, Ric76a, Ric81b, Tee80, Wag84]. programs [CPR69, Gan76, Kro72, Sny79]. progress [Ric76b]. project [Ash69, For73]. proof [Dun86b]. Propagation [Del96, Duk96, Ber75, Cyn90, Dry96]. properties [But79a, Dah79]. proportionality [Ste79]. Proposal [FW79, PS79, Pay86, PSW90, Aha85, DCHH85, DCDH87, Fel79a]. Proposed [Cod79, CCG^{+}85, CKP^{+}79, DL85b, Fel79c, Law74, Bol75, FRS77, KP79, Kni76a, Kni76b]. PROSE [Pfe87]. Provably [Dun87a, Dun87b, Dun88]. provision [Ein86]. PSE [GAA^{+}86]. pseudo [Xan86]. pseudo-ODE [Xan86]. pseudoinverse [Gol78]. pseudorandom [Ful87]. public [Smi88a]. Publication [Dun77, Ric76a, UF69]. Purdue [Mol68c, Mol69c]. purpose [CY79].

QMESH [Jon75]. QR [Rol90]. quadratic [GMSW85]. quadrature [Kah81b, Lyn68, Ric73]. qualifying [Ano71, Mol68b, Mol68c, Mol69b, Mol69c, Tra66, Tra67a, Tra67c]. quasi [Sor81]. quasi-Newton [Sor81]. question [For70]. questions [SAW85, Tra67a].

Radian [PH83b]. Radiotelescope [FK96]. Radix [CCG^{+}85]. Radix- [CCG^{+}85]. rank [HB76, Rol90]. rank- [Rol90]. rank-one [HB76]. Raphson [OR75]. rate [TK97]. ratings [TK97]. Rational [Kre95, LK96, PV83a, PV83b]. Reading [Ano73g]. Real [Ano73c, Lin81]. Realistic [LK96]. realization [Ein86]. really [Han73]. recognition [SB82]. reconsidered [Dun80]. Recursive [Bro73]. Reduction [PH83a, PH83b]. Reference [Car74]. refinement [Zie78a]. Reflections [Gil98]. Refutation [Kah91]. Regarding [WK95]. region [PP75]. register [Cyn90]. Reinterpreted [Del96, Duk96, Dry96]. related [EJ76]. relating [DL85a]. relations [Whi94]. release [GN84]. Relevant [RS97]. reliability [Boy79]. Reliable [Chi97]. Remark [Dry96, KK96]. Remarks [Ano98d, Ano98f]. Repeated [Han70]. Report [DW95, GOP^{+}79, GOP^{+}80, Smi83b]. Reports [Ano95b, Ano95c, Ano95d, Ano96a, Ano96b, Ano96c, Ano96d, Ano98a, Ano98b]. representation [EL81]. required [Par69]. Research [Ano73g, Ano73m, GOP^{+}79, GOP^{+}80, KM83, Ard80, Ric76a, RGO^{+}79, Ric81a, SAW85, Voi81]. resource [CM78]. responses [Fri72a]. Restructuring [Hop97]. Result [LK96]. results [Add81, Wit67]. reversed [Law88]. reversion [Law88]. Review [Hor74, Smi88a, Smi88b]. Reviews [Ano74a]. revision [Wag84]. Revisited [Duk96, LK96]. RObust [Kle78]. Robustness [Kah81a]. Role [Car92, DPS92]. root [Rol87]. Roots [Car74, McN98, Lin81]. ROSEPACK [Kle78]. Rounding [Fig95]. Roundoff [DM73]. Routine [HS93, Hop97, Tur69]. Routines [RW71b, Dun90c, Wit67]. rule [Dun80, Law88, Mol74]. Rules [AA96]. Runge [ES87]. Rutishauser [Zie81].

Safe [Lam79]. Saving [Bus87]. scaling [EH79]. Scandinavian [Ein78]. scheme [BHK83, Bol75, GS74]. Science [Ano71, Ano73d, Ard80, Cha87]. Scientific [Bat81, Cod71, Fos91, CG89, Dun72, Dun76a, Dun90b, GAA^{+}86, Kul83, KM83, Smi83c]. Search [Gro92, Rol97]. Second [DW95, FJ79]. section [Fri73]. Seminar [Ano73m, Kob92]. separably [Lam79]. September [Ano73i]. sequence [Pro88, Zah89]. Sequential [Ano73q, Hor74, Eva89, GMSW85]. Series [Law88, DK72]. service [Ile85]. services [PP82]. Sessions [Ng73]. Set [FKK96, DCHH85, DCDH87]. Set-Valued [FKK96]. shared [GN88]. sharp [Pri75]. shelf [CS83]. shell [CW75a]. short [Gla85]. should [Dun90c]. SIAM [Ano73n]. SIGMetrics [Ano93b]. SIGNUM [Ano75, Dry96, ACM79, Ano79, Fri72a, Han77, Rob73, Ske79b, Uso69]. SIGNUMs [Rad73]. SIGs [RS97]. simple [Tee72]. Simulations [Chi97]. sin [Tur69]. sin [?]Turner:1969:DSC. Single [PV83a, PV83b]. Single-valued [PV83a, PV83b]. singular [Ham85]. Site [Ric96]. Six [Dri82]. sizes [Cyn90]. SLATEC [VH82]. Slow [GLJ^{+}91]. Slowing [Cyn90]. soft [Kro87]. Software [ACM79, Ano73l, Ano73p, Ano75, BB91, Bri95, But79b, Car92, Cod79, Cod80, DPS92, Fel79a, LS75, Pow73, Ric91, Sch79b, Smi98, AK79, Ano79, Bat75, BHK83, Boy79, Cli82, CG79, Don84, DG85, Dri82, EJ76, Ein78, For78, FC84, Hak85, HKL73, Kel76, Kro71, Nas76, Nas90, Ric81a, Sch79c, Smi79, Wil86]. Solution [Str81, HB76, Mad75, ST75, Str77, SS75]. solve [Chi86]. solver [Chi85]. solvers [Han69, Hin76, Hin80, War76]. Solving [DZ88b, Ric96, Bra90, GAA^{+}86, Law72, Pri75, Ric76b, SW75, Fos91]. Some [Chi97, Dah79, GN88, GMSW85, Lyn68, Mac89, Ort66, Ric76a, SG72, Wis78, CPR69, Tem83]. Space [KS97, NK97, Bus87]. Sparse [DGLP82, DL85b, ES75a, GN81, GN84, GN88, Sor81]. Sparsity [DEGR88]. SPARSPAK [GN81, GN84, Lew83]. Special [Mac93, Ng73, CY79, Nas81]. specification [ABB^{+}74]. sphere [Mol69a]. spline [Bus87]. split [Cha87]. splitting [Joh75]. sponsored [Kob92]. Square [Car74, Rol87]. squares [Bra87, KCD90, Ste75]. Stability [But79a, Fuj79, Wan79]. Stabilize [Bri94]. stable [WB79]. Standard [Cod79, CCG^{+}85, CKP^{+}79, Fel79c, Kre96, PS79, Don84, KP79, Law74, PSW90, FRS77, Rei93]. standardization [ABB^{+}74, Bra79, Eng74]. Standards [RS97]. Stanford [Ano73a, Ano71, For70]. startup [Hin76]. State [Mol69d, Tra68]. Statement [Ano97]. statistical [BHK83]. Statistics [ADHR73, Ano73f, Kle78, CCHT67, Ham85, Mac89]. Status [Smi83b, Wag84]. Stefan [ST75]. Step [Ehl74, But73, See80]. Stiff [Cur73, BBG88, BL73, Ehl74, Fuj79, GS75, HB76, Lam79, Str77]. stiffness [SB82]. strategies [War76]. stress [Ber75]. Structural [Fuj79]. structure [DEGR88]. structured [CW75b, Hul73]. structures [CW75a, Sch79a]. Studies [FK96]. study [Ard80]. submatrix [Pro88]. subprograms [DCDH87, Law74, SS75, DL85a, DL85b, DCHH85]. subroutine [Boy75, Fox78, TC75, Uso69, VH82]. Subroutines [ADHR73, Bur73]. subsurface [EHHM75]. subtraction [Dun86b]. Subtractive [Del96, Duk96, Dry96]. Suggestions [Chi97]. Summary [Add81, Kob92, RW71b, Voi81, Fri72a]. Summation [Dun89c]. SuperPascal [Bri95]. Supplementary [Ano73o, DT76]. Support [Ris76]. surface [EHHM75]. Survey [Ein78, Law68, Ric81b, Fri72a, Ede91]. Swedish [Ano73p]. switches [Chi86]. Symbolic [Ste96b, dP87]. symmetric [Bra90]. Symmetry [Che97]. Symposium [Ano73c, Ano73a, Ano73k, Ano73q, Ano73r, KM83, Ano86]. syntax [Law72]. System [Ede91, Dri79, ES75b, Sny79, Tan75, CCHT67]. System/360 [CCHT67]. Systems [Fel79c, Bra90, Cyn90, Fuj79, GP87, Gre79, HB76, Kni87, Lam79, Mol74, Rhe69, Ric76a, Str76].

tabulations [Wit67]. tailoring [Boy79]. TAMPR [Dri79]. TC [Ano95a]. teach [OC70]. Teaching [BB91, Car92]. Technical [Ano74b]. technique [Jon82]. techniques [NDM75, Rup75]. Technology [RS97]. Teddington [Ano73r]. TEDDY2 [PP75]. Tee [Zie74]. term [Fri72b, Gla85]. terminal [SW75]. terminology [DT76]. Test [Dun86b, Zie78b, Add81, Cur73, DGLP82, EH77, WK95, Wit67, Zie73, Zie81]. testing [Ost79]. tests [Lyn68]. Texas [Ano73i, Ano73k]. text [Dri79, Sny79]. th [Str77]. their [TK97]. theorem [Ste75]. Theory [Ano73k]. Theses [Ano93a, Ano96e]. Thomas [KM83]. thoughts [Ort66]. Three [FRS77, Tre85]. time [Bus87]. Timing [Rol90]. TIPS [Han77]. title [Law79]. Tolerance [Ste79, Kro71]. tool [CY79, Ste80]. TOOLPACK [Mil79, Ile85]. Toolpack/1 [Ile85]. Tools [BD95, Smi83c, CG89, LE85, Ost79, Pre79]. Topics [Fra90, EJ76]. Toronto [Add81]. Transactions [Ano93c]. Transfer [Gan76, Ste91, Pay86]. Transformation [Hop97, Han73]. Transformational [Boy79]. transformations [Bra87, Rol90].

Transition [Smi98]. Transportability [Du 79]. Transportable [Fel79a, For78, Sny79]. Trap [Tyd94, Ste68]. Traub [Hor74]. treating [Rup75]. triangularization [Hua78]. trick [Xan86]. tridiagonal [Eva89]. Trigonometric [PH83a, PH83b]. Turing [HAC^{+}85]. Two [OR75, Hin80, Ker72, Par69, Smi79]. type [CW75a, DZ88a]. types [Ein86]. Typical [GS75].

UCINA [McK81]. Uncertainty [FK96]. underflow [Lin81]. Unified [Ris76]. unique [Rol88]. Univ [Ano73a, Ano73n]. Univ. [Ano73f, Ano73g, Ano73i]. Univariate [Aki96]. University [Ano71, Ano73q, Ano73m, McK81, Mol68c, Mol69c]. Unix [Dit93]. UNIX^{T M} [Sch79b, Sch79c]. Unknown [WK96]. Unknown-But-Bounded [WK96]. unstable [Cyn90]. Update [DDHH86, DDCH89]. updates [HB76, Rol90]. Use [ADHR73, Chi86, Cyn88, Dun89a, HB76, MC97, TK97]. user [Boy75, BKM88, TC75]. users [Dun76a]. Using [Cyn87, Jer89, Ost79, Vol79, Bat81, Cle84, Don84, Nas90, SW75]. utilities [Gar86].

v [Tem83]. value [ABB^{+}74, DM76, ES87, Ham85, Hin80, Wis78]. Valued [FKK96, PV83a, PV83b]. variable [War76]. variables [DZ88b]. Variants [DS94]. various [Don84]. verification [Ost79]. version [Cod88, IFI92]. via [Bai98, DG85]. virtual [Str81]. Voting [Fel73].

War [Kos96]. Waterloo [GN81, GN84]. Watson [KM83]. wave [Ber75]. wavelets [Kob92]. Web [Ric96]. weight [Dun76b]. which [Cha87]. Wisconsin [Ano73m]. Wisconsin-Madison [Ano73m]. within [Whi94]. without [CG85, ES75a]. Word [CCG^{+}85]. Word-length-independent [CCG^{+}85]. work [Wag84]. Working [Ano95a, Ein73, Fos91, Jon79, Fel79a]. Workshop [DW95, DW96, FJ79]. Worldwide [Ano96f]. worthwhile [Bra87]. Would [Hul73]. written [Dri82].

Year [Ano95e]. York [Ano73j].

Zero [Che97, OR75]. Zero-Finding [Che97]. zeros [Tra68]. Zielke [Tee74]. Zuse [Gil98].

References

Al-Alaoui:1996:CNI

[AA96] Mohamad Adnan Al-Alaoui. A class of numerical integration rules with first order derivatives. ACM SIGNUM Newsletter, 31(2):25–44, April 1996. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Aird:1974:NSV

[ABB^{+}74] Thomas J. Aird, Edward L. Battiste, Nancy E. Bosten, Henry L. Darilek, and Walton C. Gregory. Name standardization and value specification for machine dependent constants. ACM SIGNUM Newsletter, 9(4):11–13, October 1974. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

ACM:1979:PSC

[ACM79] ACM, editor. Proceedings of the SIGNUM Conference on the Programming Environment for Development of Numerical Software. ACM Press, New York, NY 10036, USA, 1979.

Addison:1981:SIT

[Add81] Clifford Addison. Summary of the Illinois–Toronto OlympiODE test results. ACM SIGNUM Newsletter, 16(4):19–22, December 1981. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Aird:1973:SUM

[ADHR73] T. Aird, D. Dodson, E. Houstis, and J. Rice. Statistics on the use of mathematical subroutines from a computer center library. ACM SIGNUM Newsletter, 8(4):8–9, October 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Aharonian:1985:MBP

[Aha85] G. Aharonian. Multi-language BLAS: a proposal. ACM SIGNUM Newsletter, 20(3):48–49, July 1985. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Aird:1979:IES

[AK79] T. J. Aird and D. G. Kainer. The IMSL environment for software development. In ACM [ACM79], pages 83–85.

Akima:1996:NLM

[Aki96] Hiroshi Akima. Note on local methods of univariate interpolation. ACM SIGNUM Newsletter, 31(2):2– 5, April 1996. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1971:SUC

[Ano71] Anonymous. Stanford University Computer Science Department Ph.D. qualifying exam: numerical analysis. ACM SIGNUM Newsletter, 6(2):11–12, May 1971. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1973:ISM

[Ano73a] Anonymous. 8th International Symposium on Mathematical Programming, Stanford Univ., August 27– 31, 1973. ACM SIGNUM Newsletter, 8(3):8, July 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1973:ANC

[Ano73b] Anonymous. ACM 73 National Conference, Atlanta, August 27– 29, 1973. ACM SIGNUM Newsletter, 8(3):8, July 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1973:ASC

[Ano73c] Anonymous. AMS symposium on the complexity of the real computational process. ACM SIGNUM Newsletter, 8(3):5, July 1973. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1973:CSC

[Ano73d] Anonymous. Computer science conference, Detroit Hilton, February 12–14, 1974. ACM SIGNUM Newsletter, 8(3):9, July 1973. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1973:CC

[Ano73e] Anonymous. Conference calendar. ACM SIGNUM Newsletter, 8(3): 3, July 1973. CODEN SNEWD6.

ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1973:CCP

[Ano73f] Anonymous. Conference on computational problems in statistics, Univ. of Essex, England, July 25– 26, 1973. ACM SIGNUM Newsletter, 8(3):7–8, July 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1973:CRM

[Ano73g] Anonymous. Conference on research in mathematics, Univ. of Reading, England, January 7–8, 1974. ACM SIGNUM Newsletter, 8(3):9, July 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1973:ICCb

[Ano73h] Anonymous. IFIP Congress 74 — call for papers. ACM SIGNUM Newsletter, 8(3):10, July 1973. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1973:ICCa

[Ano73i] Anonymous. International conference on computational methods in nonlinear mechanics, Univ. of Texas, Austin, September 23–25, 1973. ACM SIGNUM Newsletter, 8(3):9, July 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1973:NCC

[Ano73j] Anonymous. National Computer Conference, New York, June 4– 8, 1973. ACM SIGNUM Newsletter, 8(3):7, July 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1973:PPA

[Ano73k] Anonymous. Papers presented at the ACM Symposium on Theory of Computing, Austin, Texas, April 30–May 2. ACM SIGNUM Newsletter, 8(3):6, July 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1973:PPL

[Ano73l] Anonymous. Papers presented at the Loughborough Mathematical Software Conference. ACM SIGNUM Newsletter, 8(3):4–5, July 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1973:SGI

[Ano73m] Anonymous. Seminar on generalized inverses and applications, The Mathematics Research Center University of Wisconsin-Madison, October 8–10, 1973. ACM SIGNUM Newsletter, 8(3):8, July 1973. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1973:SFM

[Ano73n] Anonymous. SIAM Fall Meeting, Univ. of Iowa, Iowa City, Iowa, October 8–10, 1973. ACM SIGNUM Newsletter, 8(3):9, July 1973. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1973:SEI

[Ano73o] Anonymous. Supplementary election information. ACM SIGNUM Newsletter, 8(3):1–2, July 1973.

CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1973:SMS

[Ano73p] Anonymous. Swedish mathematical software development. ACM SIGNUM Newsletter, 8(3):11, July 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1973:SCS

[Ano73q] Anonymous. Symposium on complexity of sequential and parallel numerical algorithms, Carnegie-Mellon University, May 16-18, 1973. ACM SIGNUM Newsletter, 8(3):6– 7, July 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1973:SNM

[Ano73r] Anonymous. A symposium on numerical methods for constrained optimization, National Physical Laboratory, Teddington, England, January 10–11, 1974. ACM SIGNUM Newsletter, 8(3):9, July 1973. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1974:BR

[Ano74a] Anonymous. Book reviews. ACM SIGNUM Newsletter, 9(2):??, April 1974. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1974:TC

[Ano74b] Anonymous. Technical communications. ACM SIGNUM Newsletter, 9(2):??, April 1974. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1975:PSM

[Ano75] Anonymous. Proceedings of the SIGNUM Meeting on Software for Partial Differential Equations. ACM SIGNUM Newsletter, 10(4):??, December 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1979:PSC

[Ano79] Anonymous. Proceedings of the SIGNUM conference on the programming environment for development of numerical software. ACM SIGNUM Newsletter, 14(1):28–117, March 1979. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1986:MBM

[Ano86] Anonymous. Martin Beale memorial symposium. ACM SIGNUM Newsletter, 21(3):24, July 1986. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1991:MLA

[Ano91] Anonymous. MATH77 library available from COSMIC. ACM SIGNUM Newsletter, 26(1):5–??, January 1991. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1993:APT

[Ano93a] Anonymous. Abstracts of Ph.D. theses. ACM SIGNUM Newsletter, 28(3):2–??, July 1993. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1993:CPA

[Ano93b] Anonymous. Call for papers: 1994 ACM SIGMetrics Conference. ACM SIGNUM Newsletter, 28(3):18–??, July 1993. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1993:CPE

[Ano93c] Anonymous. Call for papers: Electronic Transactions on Numerical Analysis (ETNA). ACM SIGNUM Newsletter, 28(3):6–??, July 1993. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1995:ITW

[Ano95a] Anonymous. IFIP TC 2 Working Conference: Announcement. ACM SIGNUM Newsletter, 30(2):10–??, July 1995. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1995:IRa

[Ano95b] Anonymous. Institutional reports.

ACM SIGNUM Newsletter, 30(2):3– ??, July 1995. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1995:IRb

[Ano95c] Anonymous. Institutional reports.

ACM SIGNUM Newsletter, 30(3): 29–??, 1995. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1995:IRc

[Ano95d] Anonymous. Institutional reports.

ACM SIGNUM Newsletter, 30(4): 30–??, October 1995. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1995:ILA

[Ano95e] Anonymous. International Linear Algebra Year — announcement. ACM SIGNUM Newsletter, 30(3): 36–??, 1995. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1995:IA

[Ano95f] Anonymous. ISSAC ’95: Announcement. ACM SIGNUM Newsletter, 30(2):12–??, July 1995. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1996:IRa

[Ano96a] Anonymous. Institutional reports. ACM SIGNUM Newsletter, 31(1): 20–??, January 1996. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1996:IRb

[Ano96b] Anonymous. Institutional reports. ACM SIGNUM Newsletter, 31 (2):46–??, April 1996. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1996:IRc

[Ano96c] Anonymous. Institutional reports. ACM SIGNUM Newsletter, 31 (3):30–??, July 1996. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1996:IRd

[Ano96d] Anonymous. Institutional reports. ACM SIGNUM Newsletter, 31(4):

37–??, October 1996. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1996:PT

[Ano96e] Anonymous. Ph.D. theses. ACM SIGNUM Newsletter, 31(3):27–??, July 1996. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1996:WMC

[Ano96f] Anonymous. Worldwide Mathematica Conference. ACM SIGNUM Newsletter, 31(4):36–??, October 1996. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1997:MS

[Ano97] Anonymous. Mission statement.

ACM SIGNUM Newsletter, 32(1–2): 44–??, January/April 1997. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1998:CIR

[Ano98a] Anonymous. Current institutional reports. ACM SIGNUM Newsletter, 33(1):23–??, January 1998. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1998:DCI

[Ano98b] Anonymous. Departments: Current institutional reports. ACM SIGNUM Newsletter, 33(2):23–??, April 1998. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1998:DRB

[Ano98c] Anonymous. Departments: Recent books. ACM SIGNUM Newsletter, 33(2):29–??, April 1998. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1998:ER

[Ano98d] Anonymous. Editorial remarks. ACM SIGNUM Newsletter, 33(1): 3–??, January 1998. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1998:LC

[Ano98e] Anonymous. Letter from the Chair. ACM SIGNUM Newsletter, 33(1): 5–??, January 1998. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1998:LER

[Ano98f] Anonymous. Letters: Editorial remarks. ACM SIGNUM Newsletter, 33(2):3–??, April 1998. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Anonymous:1998:RB

[Ano98g] Anonymous. Recent books. ACM SIGNUM Newsletter, 33(1):29–??, January 1998. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Arden:1980:WCA

[Ard80] Bruce W. Arden, editor. What can be automated?: The computer science and engineering research study (COSERS), volume 3 of The MIT Press series in computer science. MIT Press, Cambridge, MA, USA, 1980. ISBN 0-262-01060-7. 934 pp. LCCN QA76 .W49.

Ashenhurst:1969:ECP

[Ash69] R. L. Ashenhurst. Evaluation and certification project at Argonne. ACM SIGNUM Newsletter, 4(3):14, October 1969. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Bailey:1998:FNM

[Bai98] David H. Bailey. Finding new mathematical identities via numerical computations. ACM SIGNUM Newsletter, 33(1):17–22, January 1998. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Battiste:1975:MSP

[Bat75] E. L. Battiste. Mathematical software patterns. ACM SIGNUM Newsletter, 10(1):17–20, January 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Battiste:1981:SCU

[Bat81] E. L. Battiste. Scientific computations using micro-computers. ACM SIGNUM Newsletter, 16(1):18–22, March 1981. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Bartholomew-Biggs:1991:AST

[BB91] M. C. Bartholomew-Biggs. Ada software for teaching modern computer arithmetic. ACM SIGNUM Newsletter, 26(3):17–26, July 1991. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Berzins:1988:SIN

[BBG88] M. Berzins, R. W. Brankin, and I. Gladwell. The stiff integrators in the NAG library. ACM SIGNUM Newsletter, 23(2):16–23, April 1988. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Bischof:1995:CAD

[BD95] Chris Bischof and Fred Dilley. A compilation of automatic differentiation tools presented at the 1995 International Convention on Industrial and Applied Mathematics, Hamburg. ACM SIGNUM Newsletter, 30(3):2–20, July 1995. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Bertholf:1975:NMN

[Ber75] Larry D. Bertholf. Numerical methods for nonlinear stress wave propagation (abstract). ACM SIGNUM Newsletter, 10(4):4, December 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Brown:1979:EPB

[BF79] W. S. Brown and S. I. Feldman. Environment parameters and basic functions for floating-point computation. In ACM [ACM79], pages 42– 45.

Brown:1983:ILA

[BFF83] W. S. Brown, K. W. Fong, and B. Ford. Issues in license agreements for numerical program libraries. ACM SIGNUM Newsletter, 18(2):21–23, April 1983. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Boisvert:1983:GCS

[BHK83] Ronald F. Boisvert, Sally E. Howe, and David K. Kahaner. The GAMS classification scheme for mathematical and statistical software. ACM SIGNUM Newsletter, 18(1):10–18, January 1983. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Brook:1988:GUG

[BKM88] Anthony Brook, David Kendrick, and Alexander Meeraus. GAMS, a user’s guide. ACM SIGNUM Newsletter, 23(3–4):10–11, July– October 1988. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Bjurel:1973:MSD

[BL73] G. Bjurel and B. Lindberg. Methods for stiff differential equations. ACM SIGNUM Newsletter, 8(1): 16–17, January 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Bolstad:1975:PCS

[Bol75] John Bolstad. A proposed classification scheme for computer program libraries. ACM SIGNUM Newsletter, 10(2–3):32–39, November 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Boyle:1975:NDU

[Boy75] James M. Boyle. Note on “The design of a user interface to large Fortran subroutine packages under IBM 360/370 OS”. ACM SIGNUM Newsletter, 10(2–3):40, November 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic). See [TC75].

Boyle:1979:ERE

[Boy79] James M. Boyle. Extending reliability (extended abstract): Transformational tailoring of abstract mathematical software. In ACM [ACM79], pages 57–60.

Brainerd:1979:CMA

[Bra79] Walt Brainerd. A “core + modules” approach to Fortran standardization. In ACM [ACM79], pages 32–33.

Branham:1987:OTW

[Bra87] Richard L. Branham, Jr. Are orthogonal transformations worthwhile for least squares problems? ACM SIGNUM Newsletter, 22(1): 14–19, January 1987. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Branham:1990:CPS

[Bra90] Richard L. Branham. A C program for solving banded, symmetric, positive definite systems. ACM SIGNUM Newsletter, 25(1):44–55, January 1990. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Branham:1995:AJD

[Bra95] Richard L. Branham, Jr. Algorithms for Julian dates. ACM SIGNUM Newsletter, 30(4):14–18, October 1995. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Briggs:1994:HSF

[Bri94] Keith Briggs. How to stabilize a fixed point. ACM SIGNUM Newsletter, 29(2):2–3, April 1994. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

BrinchHansen:1995:SS

[Bri95] Per Brinch Hansen. The Super-Pascal software. ACM SIGNUM Newsletter, 30(3):27, July 1995. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Brown:1973:RCC

[Bro73] R. Leonard Brown. Recursive calculation of corrector coefficients. ACM SIGNUM Newsletter, 8(4): 12–13, October 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Burstein:1973:SPF

[Bur73] Herman Burstein. Subroutines for probability functions. ACM SIGNUM Newsletter, 8(3):11, July 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Bush:1987:STS

[Bus87] George C. Bush. Saving time and space in spline calculations. ACM SIGNUM Newsletter, 22(1):9–10, January 1987. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Butcher:1973:SCC

[But73] J. C. Butcher. On step control criteria. ACM SIGNUM Newsletter, 8(1):15, January 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Butcher:1979:SPG

[But79a] J. C. Butcher. Stability properties for a general class of methods for ordinary differential equations. In Skeel [Ske79b], page ??

Butler:1979:PNE

[But79b] Margaret K. Butler. Portability and the National Energy Software Center. In ACM [ACM79], pages 87–89.

Carta:1974:HLR

[Car74] David G. Carta. Help!!: The lost reference: (A modified Newton method for square roots). ACM SIGNUM Newsletter, 9(4):9, October 1974. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Carroll:1992:RCS

[Car92] John Carroll. The role of computer software in numerical analysis teaching. ACM SIGNUM Newsletter, 27(2):2–31, April 1992. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Cody:1985:PRW

[CCG^{+}85] W. J. Cody, J. T. Coonen, D. M. Gay, K. Hanson, D. Hough, W. Kahan, R. Karpinski, J. Palmer, F. N. Ris, and D. Stevenson. A proposed radix- and word-length-independent standard for floating-point arithmetic. ACM SIGNUM Newsletter, 20(1):37–51, January 1985. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Clark:1967:PSF

[CCHT67] N. A. Clark, W. J. Cody, K. E.

Hillstrom, and E. A. Thieleker. Performance statistics of the Fortran IV (H) library for the IBM System/360. ACM SIGNUM Newsletter, 2(3):??, December 1967. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Coughran:1979:NLN

[CG79] Bill Coughran and Eric Grosse. New languages for numerical software. In ACM [ACM79], pages 73–75.

Chu:1985:NEE

[CG85] Eleanor Chu and Alan George. A note on estimating the error in Gaussian elimination without pivoting. ACM SIGNUM Newsletter, 20(2):2–7, April 1985. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Coughran:1989:PSC

[CG89] W. M. Coughran, Jr. and E. Grosse.

A philosophy for scientific computing tools. ACM SIGNUM Newsletter, 24(2–3):2–7, April/July 1989. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Chan:1987:OWS

[Cha87] Tony F. Chan. Obstacles which split computer science and numerical analysis. ACM SIGNUM Newsletter, 22(3):19–26, July 1987. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Chen:1997:SNI

[Che97] Tien Chi Chen. Symmetry and nonconvergence in iterative polynomial zero-finding. ACM SIGNUM Newsletter, 32(1–2):40– 43, January/April 1997. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Chikhliwala:1985:DIF

[Chi85] E. D. Chikhliwala. Design and implementation of a Fast Poisson solver for microcomputers. ACM SIGNUM Newsletter, 20(4):6–12, October 1985. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Chiang:1986:UMS

[Chi86] Yi Ling Chiang. Use of mathematical switches to solve differential equation problems. ACM SIGNUM Newsletter, 21(4):20–33, October 1986. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Chiang:1997:SSI

[Chi97] Yi-Ling F. Chiang. Some suggestions on the implementation of reliable simulations. ACM SIGNUM Newsletter, 32(1–2):11– 16, January/April 1997. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Coonen:1979:PSB

[CKP^{+}79] Jerome Coonen, William Kahan, John Palmer, Tom Pittman, and David Stevenson. A proposed standard for binary floating point arithmetic: Draft 5.11. ACM SIGNUM Newsletter, 14(3S (Special issue)): 4–12, October 1979. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Clemmesen:1984:IAI

[Cle84] Michael Clemmesen. Interval arithmetic implementations: using floating point arithmetic. ACM SIGNUM Newsletter, 19(4):2–8, October 1984. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Cline:1982:CMS

[Cli82] A. K. Cline. A course on mathematical software. ACM SIGNUM Newsletter, 17(1):19–23, March 1982. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Cuthill:1978:MRN

[CM78] Elizabeth Cuthill and Kenton Meals. MACSYMA, a resource for the Navy Laboratories. ACM SIGNUM Newsletter, 13(3): 15, September 1978. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Cody:1971:DHC

[Cod71] W. J. Cody. Desirable hardware characteristics for scientific computation. ACM SIGNUM Newsletter, 6(1):16–31, January 1971. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Cody:1979:IPI

[Cod79] W. J. Cody. Impact of the proposed IEEE floating point standard on numerical software. ACM SIGNUM Newsletter, 14(3S (Special issue)):29–30, October 1979. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Cody:1980:SBC

[Cod80] W. J. Cody. Software basics for computational mathematics. ACM SIGNUM Newsletter, 15(2):18–29, June 1980. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Cody:1988:NVM

[Cod88] W. J. Cody. New version of MACHAR available. ACM SIGNUM Newsletter, 23(3–4):9– 10, July–October 1988. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Cooper:1979:LEE

[Coo79] G. J. Cooper. Local error estimates for linear methods for ordinary differential equations. In Skeel [Ske79b], page ??

Corbett:1983:EAF

[Cor83] Robert Paul Corbett. Enhanced arithmetic for Fortran. ACM SIGNUM Newsletter, 18(1):24–28, January 1983. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Casaletto:1969:CSN

[CPR69] J. Casaletto, M. Pickett, and J. Rice. A comparison of some numerical integration programs. ACM SIGNUM Newsletter, 4(3):30–40, October 1969. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Corless:1983:SDM

[CS83] R. M. Corless and R. B. Simpson. An “off-the-shelf” dynamic mesh capability. ACM SIGNUM Newsletter, 18(2):23–24, April 1983. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Curtis:1973:SOD

[Cur73] A. R. Curtis. Stiff ordinary differential equation test problems. ACM SIGNUM Newsletter, 8(4):10–12, October 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Cavendish:1975:FEM

[CW75a] James C. Cavendish and James A.

Wixom. Finite element mesh generation for planar and shell type structures (abstract). ACM SIGNUM Newsletter, 10(4):35–37, December 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Charmonman:1975:SPF

[CW75b] S. Charmonman and J. L. Wagener. On structured programming in FORTRAN. ACM SIGNUM Newsletter, 10(1):21–23, January 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Crary:1979:APT

[CY79] F. D. Crary and J. M. Yohe. The AUGMENT precompiler as a tool for the development of special purpose arithmetic packages. In ACM [ACM79], pages 53–56.

Cynar:1987:UGE

[Cyn87] Sandra J. Cynar. Using Gaussian elimination for computation of the central difference equation coefficients. ACM SIGNUM Newsletter, 22(2):12–19, April 1987. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Cynar:1988:UFD

[Cyn88] Sandra J. Cynar. The use of forward difference equations in animation of physical phenomena. ACM SIGNUM Newsletter, 23(3–4): 4–8, July–October 1988. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Cynar:1990:SPE

[Cyn90] Sandra J. Cynar. Slowing propagation error caused by finite register sizes in unstable systems. ACM SIGNUM Newsletter, 25(3):21–29, July 1990. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dahlquist:1979:SPL

[Dah79] Germund Dahlquist. Some properties of linear multistep and oneleg methods for ordinary differential equations. In Skeel [Ske79b], page ??

deBoor:1990:EEa

[dB90a] Carl de Boor. An empty exercise. ACM SIGNUM Newsletter, 25(2):3– 7, April 1990. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

deBoor:1990:EEb

[dB90b] Carl de Boor. An empty exercise. ACM SIGNUM Newsletter, 25(4):2–6, October 1990. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dongarra:1987:PSL

[DCDH87] Jack Dongarra, Jeremy Du Croz, Iain Duff, and Sven Hammarling. A proposal for a set of level 3 basic linear algebra subprograms. ACM SIGNUM Newsletter, 22(3):2– 14, July 1987. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dongarra:1985:PES

[DCHH85] Jack J. Dongarra, Jeremy Du Croz, Sven Hammarling, and Richard J. Hanson. A proposal for an extended set of Fortran Basic Linear Algebra Subprograms. ACM SIGNUM Newsletter, 20(1):2–18, January 1985. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

deDoncker:1978:AEA

[dD78] Elise de Doncker. An adaptive extrapolation algorithm for automatic integration. ACM SIGNUM Newsletter, 13(2):12–18, June 1978. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dongarra:1989:UNL

[DDCH89] Jack Dongarra, Iain Duff, Jeremy Du Croz, and Sven Hammarling. An update notice on the level 3 BLAS. ACM SIGNUM Newsletter, 24(1): 9–10, January 1989. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dongarra:1986:UNE

[DDHH86] J. Dongarra, J. DuCroz, S. Hammarling, and R. Hanson. An update notice on the extended BLAS. ACM SIGNUM Newsletter, 21(4):2– 4, 1986. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Duff:1988:SSG

[DEGR88] I. S. Duff, A. M. Erisman, C. W. Gear, and J. K. Reid. Sparsity structure and Gaussian elimination. ACM SIGNUM Newsletter, 23(2):2– 8, April 1988. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Delaney:1996:EEP

[Del96] James A. Delaney. An example of error propagation reinterpreted as subtractive cancellation. ACM SIGNUM Newsletter, 31(1):8–13, January 1996. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic). See remark [Dry96].

Dongarra:1985:DMS

[DG85] Jack J. Dongarra and Eric Grosse. Distribution of mathematical software via electronic mail. ACM SIGNUM Newsletter, 20(3):45–47, July 1985. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Duff:1982:SMT

[DGLP82] Iain Duff, Roger Grimes, John Lewis, and Bill Poole. Sparse matrix test problems. ACM SIGNUM Newsletter, 17(2):22, June 1982. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dittmer:1993:EUC

[Dit93] Ingo Dittmer. Error in Unix commands dc and bc for multipleprecision-arithmetic. ACM SIGNUM Newsletter, 28(2):8–11, April 1993. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Danchick:1979:EMP

[DJ79] Roy Danchick and Mario L. Juncosa. On the existence of maximum polynomial degree Nordsieck–Gear (k, p) methods for all (k, p). In Skeel [Ske79b], page ??

Davis:1972:ASE

[DK72] William E. Davis and Hatem M.

Khalil. Algorithm and series expansion for the F distribution. ACM SIGNUM Newsletter, 7(3):21–23, October 1972. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dodson:1985:IRE

[DL85a] David S. Dodson and John G.

Lewis. Issues relating to extension of the Basic Linear Algebra Subprograms. ACM SIGNUM Newsletter, 20(1):19–22, January 1985. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dodson:1985:PSE

[DL85b] David S. Dodson and John G.

Lewis. Proposed sparse extensions to the Basic Linear Algebra Subprograms. ACM SIGNUM Newsletter, 20(1):22–25, January 1985. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dorr:1973:REC

[DM73] Fred W. Dorr and Cleve B. Moler. Roundoff error on the CDC 6600/ 7600 computers. ACM SIGNUM Newsletter, 8(2):24–26, April 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic). See [Kah71] for the original problem.

Daniel:1976:CAB

[DM76] James W. Daniel and Andrew J. Martin. Code available for boundary-value problems. ACM SIGNUM Newsletter, 11(1):43–44, May 1976. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dongarra:1984:PVC

[Don84] Jack J. Dongarra. Performance of various computers using standard linear equations software in a Fortran environment. ACM SIGNUM Newsletter, 19(1):23–26, January 1984. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dongarra:1992:LNA

[Don92] Jack Dongarra. LAPACK is now available. ACM SIGNUM Newsletter, 27(1):3–4, January 1992. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Douglas:1992:MMD

[Dou92] Craig C. Douglas. MGNet: a multigrid and domain decomposition network. ACM SIGNUM Newsletter, 27(4):2–8, October 1992. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

dePillis:1987:SPD

[dP87] John de Pillis. Symbolic packages in the development of numerical algorithms. ACM SIGNUM Newsletter, 22(3):15–18, July 1987. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dawson:1992:RLS

[DPS92] Jeffrey Dawson, Mary Payne, and Craig Schaffert. The role of LIA-1 in software portability. ACM SIGNUM Newsletter, 27(4):9–12, October 1992. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dritz:1979:PFP

[Dri79] Kenneth W. Dritz. Programmable formatting of program text: Experiences drawn from the TAMPR system. In ACM [ACM79], pages 71– 72.

Dritz:1982:SLE

[Dri82] Kenneth W. Dritz. Six language extensions to enhance the portability of mathematical software written in PL/I: background and justification. ACM SIGNUM Newsletter, 17(2): 24, June 1982. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Drygalla:1996:REE

[Dry96] V. Drygalla. Remark on “An example of error propagation reinterpreted as subtractive cancellation” by J. A. Delaney (SIGNUM Newsletter 1/96). ACM SIGNUM Newsletter, 31(2):24, April 1996. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic). See [Del96].

Douglas:1994:VMM

[DS94] Craig C. Douglas and Gordon Slishman. Variants of matrixmatrix multiplication for Fortran-90. ACM SIGNUM Newsletter, 29(2):4–6, April 1994. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dennis:1976:STN

[DT76] J. E. Dennis and R. A. Tapia. Supplementary terminology for nonlinear iterative methods. ACM SIGNUM Newsletter, 11(4):4–5, December 1976. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

DuCroz:1979:TPR

[Du 79] J. J. Du Croz. Transportability in practice — recent experience with the NAG Library. In ACM [ACM79], page 86.

Dukelow:1996:EEP

[Duk96] James S. Dukelow, Jr. An example of error propagation reinterpreted as subtractive cancellation — revisited. ACM SIGNUM Newsletter, 31(2):20–23, April 1996. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dunham:1972:HSC

[Dun72] Charles B. Dunham. On hardware for scientific computation. ACM SIGNUM Newsletter, 7(1):20, April 1972. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dunham:1976:BUS

[Dun76a] Charles B. Dunham. Background of users of scientific computing. ACM SIGNUM Newsletter, 11(1):44, May 1976. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dunham:1976:CAN

[Dun76b] Charles B. Dunham. Chebyshev approximation with non-continuous weight functions. ACM SIGNUM Newsletter, 11(1):44, May 1976. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dunham:1977:PNA

[Dun77] Charles B. Dunham. Publication of numerical algorithms. ACM SIGNUM Newsletter, 12(2):8, June 1977. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dunham:1980:CRR

[Dun80] Charles B. Dunham. Cramer’s rule reconsidered or equilibration desirable. ACM SIGNUM Newsletter, 15(4):9, December 1980. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dunham:1986:CPF

[Dun86a] Charles B. Dunham. The challenge of parallel function evaluation. ACM SIGNUM Newsletter, 21(4):34, October 1986. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dunham:1986:TES

[Dun86b] Charles B. Dunham. Test for (in)equality, subtraction, proof of correctness. ACM SIGNUM Newsletter, 21(3):27–30, July 1986. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dunham:1987:PMAa

[Dun87a] Charles B. Dunham. Provably monotone approximations. ACM SIGNUM Newsletter, 22(2):6–11, April 1987. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dunham:1987:PMAb

[Dun87b] Charles B. Dunham. Provably monotone approximations, II. ACM SIGNUM Newsletter, 22(3):30–31, July 1987. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dunham:1988:PMA

[Dun88] Charles B. Dunham. Provably monotone approximations, III. ACM SIGNUM Newsletter, 23(1):10, January 1988. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dunham:1989:ICA

[Dun89a] C. B. Dunham. Improvement of complex arithmetic by use of double elements. ACM SIGNUM Newsletter, 24(4):3–7, October 1989. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dunham:1989:PAH

[Dun89b] C. B. Dunham. Perturbation analysis of Horner’s method for nice cases. ACM SIGNUM Newsletter, 24(2–3):8–9, April/July 1989. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dunham:1989:S

[Dun89c] C. B. Dunham. Summation. ACM SIGNUM Newsletter, 24(1):14–15, January 1989. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dunham:1990:FPF

[Dun90a] C. B. Dunham. Feasibility of “perfect” function evaluation. ACM SIGNUM Newsletter, 25(4):25–26, October 1990. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dunham:1990:DFS

[Dun90b] Charles B. Dunham. Difficulties in fitting scientific data. ACM SIGNUM Newsletter, 25(3):15–20, July 1990. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dunham:1990:HAS

[Dun90c] Charles B. Dunham. How accurate should numerical routines be? ACM SIGNUM Newsletter, 25(2):8– 9, April 1990. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dunham:1991:ABA

[Dun91] Charles Dunham. Applications of best approximation. ACM SIGNUM Newsletter, 26(2):2–10, April 1991. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dongarra:1995:RSW

[DW95] Jack Dongarra and Jerzy Wasniewski. Report on the Second Workshop on Applied Parallel Computing, PARA95. ACM SIGNUM Newsletter, 30(4):28–??, October 1995. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dongarra:1996:DHW

[DW96] Jack Dongarra and Jerzy Wasniewski. Denmark hosts Workshop on Applied Parallel Computing (PARA96). ACM SIGNUM Newsletter, 31(3):28–??, July 1996. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dunham:1988:AEL

[DZ88a] C. Dunham and Z. Zhu. Approximation with exact Lagrangetype interpolation. ACM SIGNUM Newsletter, 23(2):14–15, April 1988. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dunham:1988:SEN

[DZ88b] Charles B. Dunham and C. Z. Zhu. Solving equations nonlinear in only one of N + 1 variables. ACM SIGNUM Newsletter, 23(3–4): 2–3, July–October 1988. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Edelman:1991:FAL

[Ede91] Alan Edelman. The First Annual Large Dense Linear System Survey. ACM SIGNUM Newsletter, 26 (4):6–12, October 1991. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Evans:1977:NCT

[EH77] D. J. Evans and M. Hatzopoulos. A note on the construction of test matrices with equal eigenvalues. ACM SIGNUM Newsletter, 12(3): 15–16, September 1977. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Evans:1979:COS

[EH79] D. J. Evans and M. Hatzopoulos. A comparison of optimal scaling and preconditioning. ACM SIGNUM Newsletter, 14(2):20–22, June 1979. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Edwards:1975:CCC

[EHHM75] L. L. Edwards, R. B. Hickman, J. K. Hobson, and T. C. Michels. CHAMP code calculations of surface and subsurface explosions (abstract). ACM SIGNUM Newsletter, 10(4):11–12, December 1975. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Ehle:1974:SCC

[Ehl74] Byron L. Ehle. Step control criteria for stiff equations. ACM SIGNUM Newsletter, 9(3):19–20, July 1974. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Einarsson:1973:IWG

[Ein73] Bo Einarsson. IFIP working group on numerical program libraries. ACM SIGNUM Newsletter, 8(3): 2, July 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Einarsson:1978:SSA

[Ein78] Bo Einarsson. Survey of Scandinavian activities in numerical software. ACM SIGNUM Newsletter, 13(3):17–19, September 1978. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Einarsson:1986:MLP

[Ein86] Bo Einarsson. Mixed language programming realization and the provision of data types. ACM SIGNUM Newsletter, 21(1–2):2–9, April 1986. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Edberg:1976:ABM

[EJ76] E. Edberg and J. Johansson. An annotated bibliography on mathematical software and related topics. ACM SIGNUM Newsletter, 11 (2):9–16, August 1976. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Eisenstat:1981:BAF

[EL81] S. C. Eisenstat and J. W. Lewis. Band arrays in Fortran: an elegant representation. ACM SIGNUM Newsletter, 16(3):9–11, September 1981. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Eggers:1979:HFP

[ELP79] Thomas W. Eggers, Judson S.

Leonard, and Mary H. Payne. Handling of floating point exceptions. In ACM [ACM79], pages 100–108.

Engel:1974:CFS

[Eng74] Frank Engel, Jr. Correspondence on Fortran standardization. ACM SIGNUM Newsletter, 9(4):9–10, October 1974. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Eisenstat:1975:EIS

[ES75a] Stanley C. Eisenstat and Andrew H.

Sherman. Efficient implementation of sparse nonsymmetric Gaussian elimination without pivoting (abstract). ACM SIGNUM Newsletter, 10(4):26–29, December 1975. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Engquist:1975:DSA

[ES75b] Björn Engquist and Tom Smedsaas.

DCG: a system for automatic code generation for hyperbolic problems (abstract). ACM SIGNUM Newsletter, 10(4):20–22, December 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Evans:1987:NRK

[ES87] D. J. Evans and B. B. Sanugi.

A nonlinear Runge–Kutta formula for initial value problems. ACM SIGNUM Newsletter, 22(3):27–30, July 1987. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Evans:1980:IMI

[Eva80] D. J. Evans. An implicit matrix inversion process. ACM SIGNUM Newsletter, 15(3):28–30, September 1980. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Evans:1989:CSP

[Eva89] D. J. Evans. A comparison of sequential and parallel elimination methods for tridiagonal matrices. ACM SIGNUM Newsletter, 24(1): 11–12, January 1989. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Farkas:1974:CNL

[Far74] I. Farkas. A conjecture in numerical linear algebra. ACM SIGNUM Newsletter, 9(4):10–11, October 1974. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Farkas:1975:E

[Far75] Imre Farkas.  Errata.  ACM

SIGNUM Newsletter, 10(1):14, January 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Farkas:1976:CNL

[Far76] I. Farkas. “A conjecture in numerical linear algebra”. ACM SIGNUM Newsletter, 11(1):47, May 1976. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Ford:1977:NLM

[FB77] Brian Ford and Janet Bentley. The NAG Library “machine”. ACM SIGNUM Newsletter, 12(4):23–24, 3, December 1977. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Fritsch:1984:ASM

[FC84] F. N. Fritsch and R. E. Carlson.

Availability of software for monotone interpolation. ACM SIGNUM Newsletter, 19(2):7, April 1984. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Feldstein:1973:VI

[Fel73] Alan Feldstein. Voting instructions.

ACM SIGNUM Newsletter, 8(3): 2, July 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Feldman:1979:CIW

[Fel79a] S. I. Feldman. Comments on the IFIP working group 2.5 proposal on “Parameterization of the Environment for Transportable Numerical Software”. ACM SIGNUM Newsletter, 14(2):18–20, June 1979. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic). See [For78].

Feldman:1979:PLE

[Fel79b] S. I. Feldman. The programming language EFL. In ACM [ACM79], pages 76–79.

Feldman:1979:IPS

[Fel79c] Stuart I. Feldman. The impact of the proposed standard for floating point arithmetic on languages and systems. ACM SIGNUM Newsletter, 14(3S (Special issue)):31–32, October 1979. CODEN SNEWD6.

ISSN 0163-5778 (print), 1558-0237 (electronic).

Figueroa:1995:WDR

[Fig95] Samuel A. Figueroa. When is double rounding innocuous? ACM SIGNUM Newsletter, 30 (3):21–26, July 1995. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic). URL

http://www.cs.nyu.edu/csweb/
Research/Theses/figueroa_sam.
pdf.

Fong:1979:NSD

[FJ79] Kirby W. Fong and Rondall E. Jones. Notes from the Second Department of Energy Library Workshop. In ACM [ACM79], pages 112– 115.

Finkelstein:1996:CSC

[FK96] Andrei M. Finkelstein and Misha Koshelev. Case studies of choosing a numerical differentiation method under uncertainty: Computer-aided design and radiotelescope network design. ACM SIGNUM Newsletter, 31(3):9–26, July 1996. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Finkelstein:1996:AEE

[FKK96] Andrei Finkelstein, Olga Kosheleva, and Vladik Kreinovich. Astrogeometry, error estimation, and other applications of set-valued analysis. ACM SIGNUM Newsletter, 31 (4):3–25, October 1996. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Forsythe:1970:SEQ

[For70] G. E. Forsythe. Stanford exam question. ACM SIGNUM Newsletter, 5(2):16, August 1970. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Ford:1973:NAG

[For73] B. Ford. The Nottingham Algorithms Group (NAG) project. ACM SIGNUM Newsletter, 8(2):16–21, April 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Ford:1978:PET

[For78] Brian Ford. Parameterization of the environment for transportable numerical software. ACM SIGNUM Newsletter, 13(2):20–23, June 1978. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic). See comments [Fel79a].

Fosdick:1991:IWC

[Fos91] Lloyd Fosdick. IFIP Working Conference on Programming Environments for High-Level Scientific Problem Solving. ACM SIGNUM Newsletter, 26(4):13–??, October 1991. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Fox:1978:PPS

[Fox78] Phyllis A. Fox. PORT: a portable subroutine library. ACM SIGNUM Newsletter, 13(1):14–15, March 1978. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Fox:1979:DPP

[Fox79] Phyllis Fox. Design principles of the PORT library. In ACM [ACM79], pages 80–82.

Francis:1981:ECN

[Fra81] M. D. Francis. Error control in the numerical inversion of the eigenvector matrix. ACM SIGNUM Newsletter, 16(1):26, March 1981. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Franke:1990:TCA

[Fra90] Richard Franke. Topics in computer-aided geometric design– 1990. ACM SIGNUM Newsletter, 25(4):7–16, October 1990. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Fritsch:1972:SNN

[Fri72a] F. N. Fritsch. SIGNUM 1971 nonacademic numerical mathematics survey: a brief summary of responses. ACM SIGNUM Newsletter, 7(1):14–18, April 1972. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Fritsch:1972:TDT

[Fri72b] F. N. Fritsch. Toward a definition of the term numerical mathematician. ACM SIGNUM Newsletter, 7(1):19, April 1972. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Fritsch:1973:PNM

[Fri73] Fred N. Fritsch. Profile on the numerical mathematics section of the Lawrence Livermore Laboratory. ACM SIGNUM Newsletter, 8 (4):9–10, October 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Ford:1977:TPA

[FRS77] B. Ford, J. K. Reid, and B. T.

Smith. Three proposed amendments to the draft proposed ANS FOR-TRAN Standard. ACM SIGNUM Newsletter, 12(1):18–20, March 1977. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Fujita:1979:SSN

[Fuj79] Shohei Fujita. Structural stability and the numerical methods for stiff differential systems. In Skeel [Ske79b], page ??

Fulda:1986:NMD

[Ful86] Joseph S. Fulda. Numerical methods disguised. ACM SIGNUM Newsletter, 21(3):31–32, July 1986. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Fullerton:1987:NPN

[Ful87] James Fullerton. Note on a pseudorandom number generator. ACM SIGNUM Newsletter, 22(4):3–5, October 1987. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Fraley:1979:PED

[FW79] Bob Fraley and Steve Walther.

Proposal to eliminate denormalized numbers. ACM SIGNUM Newsletter, 14(3S (Special issue)):22–23, October 1979. CODEN SNEWD6.

ISSN 0163-5778 (print), 1558-0237 (electronic).

Gaffney:1986:NTP

[GAA^{+}86] P. W. Gaffney, C. A. Addison, B. Andersen, S. Bjørnestad, R. E. England, P. M. Hanson, R. Pickering, and M. G. Thomason. NEXUS: towards a problem solving environment (PSE) for scientific computing. ACM SIGNUM Newsletter, 21(3):13–24, July 1986. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Gander:1976:TNP

[Gan76] W. Gander. Transfer of numerical programs. ACM SIGNUM Newsletter, 11(1):45, May 1976. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Gardner:1986:DDP

[Gar86] Timothy J. Gardner. DPUP: a distributed processing utilities package. ACM SIGNUM Newsletter, 21 (4):5–19, October 1986. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Giloi:1998:KZR

[Gil98] Wolfgang K. Giloi. Konrad Zuse: Reflections on the 80th birthday of the German computing pioneer. ACM SIGNUM Newsletter, 33(2):11–16, April 1998. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Gladwell:1985:NLO

[Gla85] I. Gladwell. The NAG library ordinary differential equations chapter and short term plans for its extension. ACM SIGNUM Newsletter, 20(3):35–40, July 1985. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Genz:1991:FGG

[GLJ^{+}91] Alan Genz, Zongli Lin, Charles Jones, Dali Luo, and Thorsten Prenzel. Fast Givens goes slow in MATLAB. ACM SIGNUM Newsletter, 26(2):11–16, April 1991. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Gill:1985:SII

[GMSW85] Philip E. Gill, Walter Murray, Michael A. Saunders, and Margaret H. Wright. Some issues in implementing a sequential quadratic programming algorithm. ACM SIGNUM Newsletter, 20(2):13–19, April 1985. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

George:1981:BDS

[GN81] Alan George and Esmond Ng. A brief description of SPARSPAK Waterloo sparse linear equations package. ACM SIGNUM Newsletter, 16(2):17–20, June 1981. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

George:1984:NRS

[GN84] Alan George and Esmond Ng. A new release of SPARSPAK: the Waterloo sparse matrix package. ACM SIGNUM Newsletter, 19(4):9–13, October 1984. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

George:1988:SSM

[GN88] Alan George and Esmond Ng. Some shared memory is desirable in parallel sparse matrix computation. ACM SIGNUM Newsletter, 23(2):9– 13, April 1988. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Goldstein:1978:FDP

[Gol78] Marvin J. Goldstein. Further decomposition of the pseudoinverse and eigensystem of a Hermitian persymmetric matrix. ACM SIGNUM Newsletter, 13(1):12–13, March 1978. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Gear:1979:NCR

[GOP^{+}79] C. W. Gear, J. M. Ortega, B. Parlett, J. R. Rice, M. Schultz, L. F. Shampine, and P. Wolfe. Numerical computation: A report on past, present, and future research. ACM SIGNUM Newsletter, 14(si-1):1–48, February 1979. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic). Also published in [GOP^{+}80].

Gear:1980:NCR

[GOP^{+}80] C. W. Gear, J. M. Ortega, B. Parlett, J. R. Rice, M. Schultz, L. F. Shampine, and P. Wolfe. Numerical computation: A report on past, present, and future research. In Arden [Ard80], pages 51–136. ISBN 0-262-01060-7. LCCN QA76 .W49. Also published in [GOP^{+}79].

Govaerts:1987:ACD

[GP87] W. Govaerts and J. D. Pryce. An alternative to Chan’s deflation for bordered systems. ACM SIGNUM Newsletter, 22(1):12–13, January 1987. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Gregory:1979:EFC

[Gre79] Robert Todd Gregory. Error-free computation with finite number systems. ACM SIGNUM Newsletter, 14(3):9–16, September 1979. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Grosse:1991:NNB

[Gro91] Eric Grosse. Netlib news: big files. ACM SIGNUM Newsletter, 26(4):4–5, October 1991. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Grosse:1992:NNS

[Gro92] Eric Grosse. Netlib news: Search for files. ACM SIGNUM Newsletter, 27(3):2–4, July 1992. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Gangi:1974:MHD

[GS74] Anthony F. Gangi and James N.

Shapiro. A modified Hotelling deflation scheme for inverting matrices. ACM SIGNUM Newsletter, 9 (1):10–11, January 1974. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Gordon:1975:TPS

[GS75] M. K. Gordon and L. F. Shampine. Typical problems for stiff differential equations. ACM SIGNUM Newsletter, 10(2–3):41, November 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Gupta:1972:AOA

[Gup72] K. K. Gupta. Applications oriented algorithms. ACM SIGNUM Newsletter, 7(2):22–23, July 1972. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Hull:1985:NT

[HAC^{+}85] T. E. Hull, A. Abrham, M. S. Cohen, A. F. X. Curley, C. B. Hall, D. A. Penny, and J. T. M. Sawchuk. Numerical Turing. ACM SIGNUM Newsletter, 20(3):26–34, July 1985. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Hake:1985:MSK

[Hak85] J. Fr. Hake. Mathematical software at KFA. ACM SIGNUM Newsletter, 20(2):20–30, April 1985. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Hammarling:1985:SVD

[Ham85] Sven Hammarling. The singular value decomposition in multivariate statistics. ACM SIGNUM Newsletter, 20(3):2–25, July 1985. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Hansen:1968:IEE

[Han68] Eldon Hansen. An illustrative example in error analysis. ACM SIGNUM Newsletter, 3(3):??, July 1968. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Hanson:1969:CLE

[Han69] Richard J. Hanson. Certifying linear equation solvers. ACM SIGNUM Newsletter, 4(3):21–29, October 1969. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Hansen:1970:RA

[Han70] Eldon Hansen. Repeated additions.

ACM SIGNUM Newsletter, 5(1):11– 12, March 1970. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Hanson:1973:FGT

[Han73] Richard J. Hanson. Is the fast Givens transformation really fast? ACM SIGNUM Newsletter, 8(4):7, October 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Hanson:1977:TOL

[Han77] Richard J. Hanson. TIPS for organizing a local ACM-SIGNUM group. ACM SIGNUM Newsletter, 12(4):20, December 1977. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Hindmarsh:1976:URO

[HB76] A. C. Hindmarsh and G. D. Byrne.

On the use of rank-one updates in the solution of stiff systems of ordinary differential equations. ACM SIGNUM Newsletter, 11(3):23–27, October 1976. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Higginbotham:1986:AF

[Hig86] T. F. Higginbotham. Another factor of (10^{31} − 1)/9. ACM SIGNUM Newsletter, 21(3):12, July 1986. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Hindmarsh:1976:NSO

[Hin76] A. C. Hindmarsh. Note on the startup of ODE solvers. ACM SIGNUM Newsletter, 11(4):12–13, December 1976. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Hindmarsh:1980:LLT

[Hin80] Alan C. Hindmarsh. LSODE and LSODI, two new initial value ordinary differential equation solvers. ACM SIGNUM Newsletter, 15(4): 10–11, December 1980. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Hanson:1973:IEP

[HKL73] R. J. Hanson, F. T. Krogh, and C. L. Lawson. Improving the efficiency of portable software for linear algebra. ACM SIGNUM Newsletter, 8(4):16, October 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Hodge:1995:AFD

[Hod95] H. K. Hodge. Achieving a full deck. ACM SIGNUM Newsletter, 30 (4):19–26, October 1995. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Hodge:1996:MI

[Hod96] H. K. Hodge. More interesting. ACM SIGNUM Newsletter, 31(1):18, January 1996. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Hopkins:1997:RBL

[Hop97] Tim Hopkins. Restructuring the BLAS Level 1 routine for computing the modified Givens transformation. ACM SIGNUM Newsletter, 32 (4):2–14, October 1997. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Horowitz:1974:RCS

[Hor74] Ellis Horowitz. Review of “Complexity of Sequential and Parallel Numerical Algorithms by J. F. Traub”. ACM SIGNUM Newsletter, 9(2):20–21, April 1974. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Hopkins:1993:CEM

[HS93] Tim Hopkins and John Slater. A comment on the Eispack machine epsilon routine. ACM SIGNUM Newsletter, 28(4):2–6, October 1993. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Huang:1976:NPC

[Hua76] C. P. Huang. Note on the preconditioning of matrices. ACM SIGNUM Newsletter, 11(4):11–12, December 1976. CODEN SNEWD6.

ISSN 0163-5778 (print), 1558-0237 (electronic).

Huang:1978:TAM

[Hua78] C. P. Huang. On the triangularization of arbitrary matrices. ACM SIGNUM Newsletter, 13(3): 16–17, September 1978. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Hull:1973:WYB

[Hul73] T. E. Hull. Would you believe structured Fortran? ACM SIGNUM Newsletter, 8(4):13–16, October 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Hull:1979:DFP

[Hul79] T. E. Hull. Desirable floating-point arithmetic and elementary functions for numerical computation. In ACM [ACM79], pages 96–99.

IFIF:1992:CVD

[IFI92] IFIP Working Group 2.5 (Numerical Software). Comments on version 3.1 of draft ISO/IEC 10967:1991 Language Compatible Arithmetic. ACM SIGNUM Newsletter, 27(1): 2–3, January 1992. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

IFIP:1993:ECE

[IFI93] IFIP Working Group 2.5. The enable construct for exception handling in Fortran 90. ACM SIGNUM Newsletter, 28(4):7–16, October 1993. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Iles:1985:TDS

[Ile85] R. M. J. Iles. Toolpack/1 distribution service. ACM SIGNUM Newsletter, 20(3):41–44, July 1985. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Jensen:1971:MP

[Jen71] Paul S. Jensen. A matrix problem.

ACM SIGNUM Newsletter, 6(2): 10, May 1971. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Jerrell:1989:ADU

[Jer89] Max E. Jerrell. Automatic differentiation using almost any language. ACM SIGNUM Newsletter, 24(1):2–9, January 1989. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Johnson:1975:SCA

[Joh75] W. E. Johnson. The splitting concept applied to the OIL family of codes (abstract). ACM SIGNUM Newsletter, 10(4):23, December 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Jones:1975:QMG

[Jon75] Rondall E. Jones. The QMESH mesh generation package (abstract). ACM SIGNUM Newsletter, 10(4): 31–34, December 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Jones:1979:ADA

[Jon79] Rondall E. Jones. Activities of the DOE Advanced Computing Committee Language Working Group. In ACM [ACM79], pages 116–117.

Jones:1982:NAD

[Jon82] Bush Jones. A note on Aitken’s δ^{2} technique. ACM SIGNUM Newsletter, 17(2):23, June 1982. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Kahan:1968:FAC

[Kah68] W. Kahan. Four aphorisms concerning floating point hardware design. ACM SIGNUM Newsletter, 3(2): ??, July 1968. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Kahan:1971:P

[Kah71] W. Kahan. A problem. ACM SIGNUM Newsletter, 6(3):6, November 1971. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic). See response [DM73].

Kahan:1972:MHK

[Kah72] W. Kahan. In memoriam: Hirondo Kuki: Apr. 25, 1925–Dec. 28, 1971. ACM SIGNUM Newsletter, 7(1):8– 10, April 1972. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Kahan:1981:PEC

[Kah81a] W. Kahan. The programming environment’s contribution to program robustness. ACM SIGNUM Newsletter, 16(4):10–??, October 1981. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Kahaner:1981:BNQ

[Kah81b] David K. Kahaner. BASIC for numerical quadrature. ACM SIGNUM Newsletter, 16(2):21–24, June 1981. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Kahan:1991:ARL

[Kah91] W. Kahan. Analysis and refutation of the LCAS. ACM SIGNUM Newsletter, 26(3):2–15, July 1991. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Karp:1985:PPC

[Kar85] A. Karp. A parallel processing challenge. ACM SIGNUM Newsletter, 20(4):12–13, October 1985. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Khan:1990:NCN

[KCD90] Zafar Khan, Sudhir Chawla, and Dinesh Dave. A note on comparison of nonlinear least squares algorithms. ACM SIGNUM Newsletter, 25(2):10–18, April 1990. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Kelman:1976:FLA

[Kel76] Robert B. Kelman. Factors limiting academic participation in the development of large mathematical software packages. ACM SIGNUM Newsletter, 11(1):43, May 1976. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Kershaw:1972:EFT

[Ker72] D. Kershaw. The explicit factorization of two commonly occurring matrices. ACM SIGNUM Newsletter, 7(1):13, April 1972. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Koshelev:1996:WMI

[KK96] Misha Koshelev and Vladik Kreinovich. Why monotonicity in interval computations? A remark. ACM SIGNUM Newsletter, 31(3):4–8, July 1996. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Klema:1978:RRS

[Kle78] Virginia Klema. ROSEPACK: RObust Statistics Estimation PACKage. ACM SIGNUM Newsletter, 13(2):18–19, June 1978. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Kulisch:1983:NAS

[KM83] U. Kulisch and W. Miranker, editors. A new approach to scientific computation: Proceedings of the Symposium on a New Approach to Scientific Computation (1982: IBM Thomas J. Watson Research Center). Academic Press, New York, NY, USA, 1983. ISBN 0-12-428660-7. LCCN QA297 .S847 1982.

Knight:1972:PJ

[Kni72] William Knight. A problem by Jensen. ACM SIGNUM Newsletter, 7(1):11–12, April 1972. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Knight:1976:DPA

[Kni76a] W. R. Knight. Draft proposed ANS FORTRAN. ACM SIGNUM Newsletter, 11(1):41–43, May 1976. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Knight:1976:LEC

[Kni76b] William R. Knight. A letter to the editor concerning draft proposed ANS FORTRAN. ACM SIGNUM Newsletter, 11(1):41–43, May 1976. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Knight:1977:DOE

[Kni77] William R. Knight. A disquieting output error. ACM SIGNUM Newsletter, 12(3):15, September 1977. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Knight:1987:AAB

[Kni87] William Knight. Another alternative for bordered systems. ACM SIGNUM Newsletter, 22(4):12, October 1987. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Kobayashi:1992:SSW

[Kob92] Mei Kobayashi. Summary of the seminar on wavelets sponsored by AVIRG. ACM SIGNUM Newsletter, 27(1):5–8, January 1992. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Koshelev:1996:CCP

[Kos96] Misha Koshelev. Computationally complicated problems of numerical computations (in particular, interval computations), biofeedback, computer games, and Gulf War: An idea. ACM SIGNUM Newsletter, 31(1):2–7, January 1996. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Kahan:1979:PFP

[KP79] W. Kahan and J. Palmer. On a proposed floating-point standard. ACM SIGNUM Newsletter, 14(3S (Special issue)):13–21, October 1979. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Kreinovich:1995:IRA

[Kre95] Vladik Kreinovich. Interval rational = algebraic. ACM SIGNUM Newsletter, 30(4):2–13, October 1995. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic). See remark [LK96].

Kreinovich:1996:WAC

[Kre96] Vladik Kreinovich. With what accuracy can we measure masses if we have an (approximately known) mass standard. ACM SIGNUM Newsletter, 31(4):26–34, October 1996. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Krogh:1971:PTE

[Kro71] Fred T. Krogh. A plea for tolerance in the evaluation of numerical methods and mathematical software. ACM SIGNUM Newsletter, 6 (3):7–8, November 1971. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Krogh:1972:OMC

[Kro72] Fred T. Krogh. Opinions on matters connected with the evaluation of programs and methods for integrating ordinary differential equations. ACM SIGNUM Newsletter, 7(3):27–48, October 1972. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Krogh:1987:MHC

[Kro87] Fred T. Krogh. Making hard constraints soft. ACM SIGNUM Newsletter, 22(1):11, January 1987. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Koshelev:1997:ESN

[KS97] Misha Koshelev and Scott Starks.

Energy from space: A new potential application of interval computations. ACM SIGNUM Newsletter, 32(3):9–13, July 1997. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Kulisch:1983:FES

[Kul83] U. Kulisch. A FORTRAN extension for scientific computation. In Kulisch and Miranker [KM83], pages 199–223. ISBN 0-12-428660-7. LCCN QA297 .S847 1982.

Lambert:1979:SPM

[Lam79] J. D. Lambert. Safe point methods for separably stiff systems. In Skeel [Ske79b], page ??

Lawson:1968:SCM

[Law68] Charles L. Lawson. Survey of computer methods for fitting curves to discrete data or approximating continuous functions. ACM SIGNUM Newsletter, 3(3):??, July 1968. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Lawson:1972:SSL

[Law72] C. L. Lawson. A syntax for solving linear equations. ACM SIGNUM Newsletter, 7(3):14–15, October 1972. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Lawson:1973:E

[Law73] C. L. Lawson. From the editor. ACM SIGNUM Newsletter, 8(3): 1, July 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Lawson:1974:PSS

[Law74] C. L. Lawson. Proposed standard subprograms for basic linear algebraic operations. ACM SIGNUM Newsletter, 9(2):21–22, April 1974. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Lawson:1979:NT

[Law79] Charles L. Lawson. No title. In ACM [ACM79], pages 38–41.

Lawson:1988:SRR

[Law88] Charles L. Lawson. Series reversion as the reversed chain rule. ACM SIGNUM Newsletter, 23(1): 7–9, January 1988. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Lipps:1985:PTL

[LE85] H. Lipps and G. A. Erskine. Programming tools for linear algebra. ACM SIGNUM Newsletter, 20(3):50–53, July 1985. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Lewis:1983:NES

[Lew83] John Gregg Lewis. Numerical experiments with SPARSPAK. ACM SIGNUM Newsletter, 18(3):12–22, July 1983. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Linnainmaa:1981:CEU

[Lin81] Seppo Linnainmaa. Combatting the effects of underflow and overflow in determining real roots of polynomials. ACM SIGNUM Newsletter, 16(2):11–16, June 1981. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Lakeyev:1996:IRA

[LK96] Anatoly V. Lakeyev and Vladik Kreinovich. “Interval Rational = Algebraic” revisited: A more computer realistic result. ACM SIGNUM Newsletter, 31(1):14–17, January 1996. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic). See [Kre95].

Lomax:1975:SCP

[LS75] H. Lomax and K. Stevens. Software for computers with parallel or pipeline architecture (abstract). ACM SIGNUM Newsletter, 10(4):25, December 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Lyness:1968:SIQ

[Lyn68] J. N. Lyness. Some interesting quadrature tests. ACM SIGNUM Newsletter, 3(2):??, July 1968. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Macleod:1989:SSG

[Mac89] A. J. Macleod. Some statistics on Gaussian elimination with partial pivoting. ACM SIGNUM Newsletter, 24(2–3):10–14, April/July 1989. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

MacLeod:1993:IPC

[Mac93] Allan J. MacLeod. An instability problem in Chebyshev expansions for special functions. ACM SIGNUM Newsletter, 28(2):2–7, April 1993. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Madsen:1975:MLN

[Mad75] Niel K. Madsen. The method of lines for the numerical solution of partial differential equations (abstract). ACM SIGNUM Newsletter, 10(4):5–7, December 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Maerten:1997:IPA

[MC97] Bart Maerten and Ronald Cools. An interactive program to approximate double integrals: an easy to use interface for Cubpack++. ACM SIGNUM Newsletter, 32(3):2– 8, July 1997. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

McKee:1981:UCI

[McK81] S. McKee. University Consortium for Industrial Numerical Analysis (UCINA). ACM SIGNUM Newsletter, 16(4):16–19, December 1981. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

McNamee:1998:CMA

[McN98] John Michael McNamee. A comparison of methods for accelerating convergence of Newton’s method for multiple polynomial roots. ACM SIGNUM Newsletter, 33(2):17–22, April 1998. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Miller:1979:CT

[Mil79] Webb Miller. A case for TOOL-PACK. In ACM [ACM79], pages 68–70.

Moler:1967:P

[ML67] C. B. Moler and J. N. Lyness. Problems. ACM SIGNUM Newsletter, 1 (3):12–13, February 1967. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Matuska:1975:AHC

[MND75] Daniel A. Matuska, Charles E.

Needham, and Richard E. Durrett. The AFWL HULL code-codes for large problems in fluid dynamics. ACM SIGNUM Newsletter, 10(4): 15–17, December 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Moler:1968:EPD

[Mol68a] C. B. Moler. Excerpts from Ph.D.

comprehensive examination. ACM SIGNUM Newsletter, 3(3):??, July 1968. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Moler:1968:PDQ

[Mol68b] Cleve Moler. Ph.D. qualifying examination: numerical analysis.

ACM SIGNUM Newsletter, 3(1):??, April 1968. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Moler:1968:PUN

[Mol68c] Cleve Moler. Purdue University: numerical analysis qualifying examination. ACM SIGNUM Newsletter, 3(2):??, July 1968. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Moler:1969:MSC

[Mol69a] C. B. Moler. More on the sphere in the corner. ACM SIGNUM Newsletter, 4(1):7, January 1969. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic). See [Ste68].

Moler:1969:PDQ

[Mol69b] C. B. Moler. Ph.D. qualifying examination: numerical analysis. ACM SIGNUM Newsletter, 4(2):13– 16, June 1969. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Moler:1969:PUN

[Mol69c] C. B. Moler. Purdue University: numerical analysis qualifying examination. ACM SIGNUM Newsletter, 4 (1):11–13, January 1969. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Moler:1969:SAM

[Mol69d] Cleve B. Moler. State of the art in matrix computations. ACM SIGNUM Newsletter, 4(1):22–28, January 1969. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Moler:1972:MGF

[Mol72] Cleve Moler. A memory of George Forsythe. ACM SIGNUM Newsletter, 7(3):8–9, October 1972. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Moler:1974:CRS

[Mol74] Cleve Moler. Cramer’s rule on 2-by-2 systems. ACM SIGNUM Newsletter, 9(4):13–14, October 1974. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Moore:1980:MIA

[Moo80] R. E. Moore. Microprogrammed interval arithmetic. ACM SIGNUM Newsletter, 15(2):30, June 1980. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Morris:1969:CAC

[Mor69] William L. Morris. On counting nary compositions. ACM SIGNUM Newsletter, 4(2):7–8, June 1969. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

More:1980:O

[MTW80] Jorge J. More, Richard A. Tapia, and Margaret H. Wright. Optimization. ACM SIGNUM Newsletter, 15 (4):16–26, December 1980. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Nash:1976:NSM

[Nas76] J. C. Nash. Numerical software for minicomputers. ACM SIGNUM Newsletter, 11(1):40–41, May 1976. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Nash:1981:FMS

[Nas81] J. C. Nash. Function minimization on special computers. ACM SIGNUM Newsletter, 16(1):14–17, March 1981. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Nash:1990:HEU

[Nas90] John C. Nash. A hypertext environment for using mathematical software: part 1 — motivations and design objectives. ACM SIGNUM Newsletter, 25(3):2–14, July 1990. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Needham:1975:AHD

[NDM75] Charles E. Needham, Richard E. Durrett, and Daniel A. Matuska. AFWL hydrocode development parallel techniques: fluid dynamics (abstract). ACM SIGNUM Newsletter, 10(4):13–14, December 1975. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Ng:1973:ASS

[Ng73] Edward W. Ng. AMS sessions on special functions: Evanston, IL, April 27–28, 1973. ACM SIGNUM Newsletter, 8(3):5, July 1973. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Nikolai:1972:FEC

[Nik72] Paul J. Nikolai. Fast eigenvalue calculations on the CDC 6000– 7000 computers. ACM SIGNUM Newsletter, 7(3):16–18, October 1972. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Nguyen:1997:ICS

[NK97] Hung T. Nguyen and Vladik Kreinovich. Intelligent control in space exploration: Interval computations are needed. ACM SIGNUM Newsletter, 32(1–2):17– 36, January/April 1997. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Oldehoeft:1970:IEP

[OC70] A. E. Oldehoeft and S. D. Conte.

Initial experiences with a program to teach numerical methods. ACM SIGNUM Newsletter, 5(2):17–20, August 1970. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Oster:1975:TDN

[OR75] C. A. Oster and R. L. Richardson. Two dimensional Newton– Raphson and a zero Jacobian. ACM SIGNUM Newsletter, 10(2–3):41– 42, November 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Ortega:1966:STN

[Ort66] J. Ortega. Some thoughts on numerical analysis. ACM SIGNUM Newsletter, 1(1):??, June 1966. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Osterweil:1979:UTV

[Ost79] Leon J. Osterweil. Using tools for verification, documentation and testing. In ACM [ACM79], pages 63–67.

Parlett:1967:EA

[Par67] B. N. Parlett. Eigenvalues, anyone? ACM SIGNUM Newsletter, 1(3):7–8, February 1967. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Parlett:1969:NNM

[Par69] B. N. Parlett. Notes on the number of multiplications required to form the product of two N × N matrices. ACM SIGNUM Newsletter, 4(2):9– 12, June 1969. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Payne:1986:PTF

[Pay86] Mary H. Payne. Proposal for the transfer of floating point data. ACM SIGNUM Newsletter, 21(3):25–26, July 1986. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Pfeiffer:1987:ADP

[Pfe87] F. W. Pfeiffer. Automatic differentiation in PROSE. ACM SIGNUM Newsletter, 22(1):2–8, January 1987. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Payne:1983:DRT

[PH83a] Mary H. Payne and Robert N. Hanek. Degree reduction for trigonometric functions. ACM SIGNUM Newsletter, 18(2):18–19, April 1983. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Payne:1983:RRT

[PH83b] Mary H. Payne and Robert N.

Hanek. Radian reduction for trigonometric functions. ACM SIGNUM Newsletter, 18(1):19–24, January 1983. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Powell:1973:MSC

[Pow73] M. J. D. Powell. The mathematical software conference at Loughborough. ACM SIGNUM Newsletter, 8(3):4, July 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Polak:1975:TPP

[PP75] S. J. Polak and N. V. Philips.

TEDDY2: a program package for 2D parabolic composite region problems (abstract). ACM SIGNUM Newsletter, 10(4):40–42, December 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Phillips:1982:NMB

[PP82] J. L. Phillips and W. G. Poole.

Numerical mathematics at Boeing computer services. ACM SIGNUM Newsletter, 17(3):21–22, September 1982. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Presser:1979:FT

[Pre79] Leon Presser. Fortran tools. In ACM [ACM79], pages 61–62.

Price:1975:AFE

[Pri75] Harvey S. Price. Application of finite element methods to solving immiscible displacement problems with sharp fronts (abstract). ACM SIGNUM Newsletter, 10(4):24, December 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Proskurowski:1988:DSM

[Pro88] Wlodzimierz Proskurowski. Decomposing a sequence of matrices that differ only in one submatrix. ACM SIGNUM Newsletter, 23(1):12–17, January 1988. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Payne:1979:DPB

[PS79] Mary Payne and William Strecker. Draft proposal for a binary normalized floating point standard. ACM SIGNUM Newsletter, 14(3S (Special issue)):24–28, October 1979. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Payne:1990:PLC

[PSW90] M. Payne, C. Schaffert, and B. Wichmann. Proposal for a language compatible arithmetic standard. ACM SIGNUM Newsletter, 25(1):2–43, January 1990. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Pyzalski:1983:SCD

[PV83a] R. Pyzalski and M. Vala. Singlevalued conversion of decimal numbers to rational fractions. ACM SIGNUM Newsletter, 18(4):22, October 1983. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Pyzalski:1983:SVC

[PV83b] R. Pyzalski and M. Vala. Singlevalued conversion of decimal numbers to rational fractions. ACM SIGNUM Newsletter, 18(4):22, October 1983. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Radbill:1973:LS

[Rad73] John Radbill. Local SIGNUMs.

ACM SIGNUM Newsletter, 8(3): 2, July 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Ris:1993:WFP

[RBSF93] Fred Ris, Ed Barkmeyer, Craig Schaffert, and Peter Farkas. When floating-point addition isn’t commutative. ACM SIGNUM Newsletter, 28(1):8–13, January 1993. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Reid:1979:FMF

[Rei79a] John Reid. Functions for manipulating floating-point numbers. ACM SIGNUM Newsletter, 14(4):11–13, December 1979. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Reinsch:1979:PPC

[Rei79b] Christian H. Reinsch. Principles and preferences for computer arithmetic. ACM SIGNUM Newsletter, 14(1):12–27, March 1979. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Reid:1980:CDP

[Rei80] John Reid. Complex double precision in association with Fortran 77. ACM SIGNUM Newsletter, 15 (1):16–17, March 1980. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Reid:1993:LIA

[Rei93] John Reid. Language Independent Arithmetic (LIA) — A Draft International Standard (DIS). ACM SIGNUM Newsletter, 28(1):2–7, January 1993. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Rice:1979:NCN

[RGO^{+}79] J. R. Rice, C. W. Gear, J. Ortega, B. Parlett, M. Schultz, L. F. Shampine, P. Wolfe, and J. F. Traub. Numerical computation: its nature and research directions. ACM SIGNUM Newsletter, 14 (3S (Special issue)):1–48, February 1979. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Rheinboldt:1969:NSE

[Rhe69] Werner C. Rheinboldt. Nonlinear systems of equations. ACM SIGNUM Newsletter, 4(1):15–21, January 1969. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Rice:1973:EAQ

[Ric73] John R. Rice. An educational adaptive quadrature algorithm. ACM SIGNUM Newsletter, 8(2):27–41, April 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic). See errata [Rob73].

Rice:1976:SDO

[Ric76a] J. R. Rice. Some data and observations on research publication in the areas of numerical computation and programming languages and systems. ACM SIGNUM Newsletter, 11(3):28–32, October 1976. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Rice:1976:APS

[Ric76b] John R. Rice. Algorithmic progress in solving partial differential equations. ACM SIGNUM Newsletter, 11(4):6–10, December 1976. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Rice:1979:PEA

[Ric79] John R. Rice. “programming effort”

analysis of the ELLPACK language. In ACM [ACM79], pages 109–111.

Rice:1981:MSL

[Ric81a] John R. Rice. Is mathematical software a legitimate research area? ACM SIGNUM Newsletter, 16(1):23–25, March 1981. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Rice:1981:SPL

[Ric81b] John R. Rice. Survey on programming language facilities for numerical computation. ACM SIGNUM Newsletter, 16(3):14–20, September 1981. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Rice:1991:PEN

[Ric91] John R. Rice. Performance evaluation of numerical software. ACM SIGNUM Newsletter, 26(1):7–??, January 1991. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Rice:1996:PSE

[Ric96] John Rice. Problem solving environments — Web site. ACM SIGNUM Newsletter, 31(4):35–??, October 1996. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Ris:1976:UDF

[Ris76] Frederic N. Ris. A unified decimal floating-point architecture for the support of high-level languages. ACM SIGNUM Newsletter, 11(3): 18–23, October 1976. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Roberts:1973:E

[Rob73] O. P. Roberts. Errata: SIGNUM Newsletter 8(2) 32 April (1973). ACM SIGNUM Newsletter, 8(3): 11, July 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic). See [Ric73].

Rolfe:1987:FIS

[Rol87] Timothy J. Rolfe. On a fast integer square root algorithm. ACM SIGNUM Newsletter, 22(4):6–11, October 1987. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Rolfe:1988:GPN

[Rol88] T. J. Rolfe. Generations of permutations with non-unique elements. ACM SIGNUM Newsletter, 23(2):24–28, April 1988. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Rolfe:1990:TCH

[Rol90] Timothy J. Rolfe. Timing comparisons of the Householder QR transformations with rank-1 and rank-2 updates. ACM SIGNUM Newsletter, 25(4):19–24, October 1990. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Rolfe:1997:ADC

[Rol97] Timothy J. Rolfe. Analytic derivation of comparisons in binary search. ACM SIGNUM Newsletter, 32(4):15–19, October 1997. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Rada:1997:NSE

[RS97] Roy Rada and James Schoening.

New standards for educational technology relevant to multiple ACM SIGs. ACM SIGNUM Newsletter, 32(1–2):37–39, January/April 1997. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Ruppel:1975:RDN

[Rup75] Hans M. Ruppel. Recent developments in numerical techniques for treating material interfaces (abstract). ACM SIGNUM Newsletter, 10(4):8–10, December 1975. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Redish:1971:EEN

[RW71a] K. A. Redish and W. A. Ward. Environment enquiries for numerical analysis. ACM SIGNUM Newsletter, 6(1):10–15, January 1971. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Redish:1971:SEE

[RW71b] K. A. Redish and W. A. Ward. Summary of “Eigenvalue & Eigenvector Routines in the McMaster Program Library”. ACM SIGNUM Newsletter, 6(2):8–9, May 1971. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Solomon:1985:MBP

[SAW85] A. D. Solomon, V. Alexiades, and D. G. Wilson. Moving boundary problems in phase change models current research questions. ACM SIGNUM Newsletter, 20(2):8–12, April 1985. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Saylor:1980:NML

[Say80] Paul E. Saylor. Numerical methods, large codes, and programmers. ACM SIGNUM Newsletter, 15(2):16–17, June 1980. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Shampine:1982:ARN

[SB82] L. F. Shampine and L. S. Baca. Automatic recognition of non-stiffness. ACM SIGNUM Newsletter, 17 (2):20–21, June 1982. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Schenk:1979:NPD

[Sch79a] Werner Schenk. Numerical precision and data structures. In ACM [ACM79], pages 34–37.

Schryer:1979:UEPa

[Sch79b] N. L. Schryer. UNIX^{T M} as an environment for producing numerical software. ACM SIGNUM Newsletter, 14(1):49–52, 1979. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Schryer:1979:UEPb

[Sch79c] N. L. Schryer. UNIX^{T M} as an environment for producing numerical software. In ACM [ACM79], pages 49–52.

Seehusen:1980:AGS

[See80] J. Seehusen. Algorithm for generating a step matrix by permutation. ACM SIGNUM Newsletter, 15 (4):12–15, December 1980. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Shampine:1972:SNE

[SG72] L. F. Shampine and M. K. Gordon.

Some numerical experiments with DIFSUB. ACM SIGNUM Newsletter, 7(3):24–26, October 1972. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Skeel:1979:ILB

[Ske79a] R. D. Skeel. An incomplete list of better FORTRAN codes for IVPS in ODES. ACM SIGNUM Newsletter, 14(2):22–24, June 1979. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Skeel:1979:SMN

[Ske79b] R. D. Skeel, editor. SIGNUM Meeting on Numerical Ordinary Differential Equations, volume 963 of Department of Computer Science Report. University of Illinois, Champaign, IL, USA, 1979.

Smith:1979:CTR

[Smi79] Brian T. Smith. A comparison of two recent approaches to machine parameterization for mathematical software. In ACM [ACM79], pages 46–48.

Smith:1983:EA

[Smi83a] B. T. Smith. EISPACK 3: announcement of availability. ACM SIGNUM Newsletter, 18(4):18, October 1983. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Smith:1983:SRF

[Smi83b] B. T. Smith. Status report on FOR-TRAN 8x as of August 1983. ACM SIGNUM Newsletter, 18(4):19–21, October 1983. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Smith:1983:TML

[Smi83c] B. T. Smith. Tools, methods and languages for scientific and engineering computation. ACM SIGNUM Newsletter, 18(4):9–11, October 1983. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Smith:1988:FPR

[Smi88a] Brian T. Smith. Fortran 8x — its public review. ACM SIGNUM Newsletter, 23(1):2–6, January 1988. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Smith:1988:RAF

[Smi88b] Brian T. Smith. A review and analysis of Fortran 8x. ACM SIGNUM Newsletter, 23(2):29–60, April 1988. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Smith:1998:TNS

[Smi98] Barry F. Smith. The transition of numerical software: From nutsand-bolts to abstraction. ACM SIGNUM Newsletter, 33(1):7–15, January 1998. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Snyder:1979:TSM

[Sny79] W. V. Snyder. A transportable system for management and exchange of programs and other text. In ACM [ACM79], pages 90–92.

Sorensen:1981:ECQ

[Sor81] Danny C. Sorensen. An example concerning quasi-Newton estimation of a sparse Hessian. ACM SIGNUM Newsletter, 16(2):8–10, June 1981. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Swarztrauber:1975:EFS

[SS75] Paul Swarztrauber and Roland Sweet. Efficient FORTRAN subprograms for the solution of elliptic partial differential equations (abstract). ACM SIGNUM Newsletter, 10(4):30, December 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Shampine:1975:EMN

[ST75] L. F. Shampine and R. J. Thompson. An efficient method for the numerical solution of one-phase Stefan problems (abstract). ACM SIGNUM Newsletter, 10(4):18–19, December 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Stenger:1968:TC

[Ste68] Frank Stenger. A trap in computations. ACM SIGNUM Newsletter, 3(3):2, July 1968. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic). See [Mol69a].

Stewart:1975:IPT

[Ste75] G. W. Stewart. An inverse perturbation theorem for the linear least squares problem. ACM SIGNUM Newsletter, 10(2–3):39–40, November 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Stetter:1979:TPO

[Ste79] Hans J. Stetter. Tolerance proportionality in ODE-codes. In Skeel [Ske79b], page ??

Stewart:1980:MTN

[Ste80] Kris Stewart. The microcomputer as a tool in numerical analysis. ACM SIGNUM Newsletter, 15(3): 27, September 1980. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Stewart:1981:MCH

[Ste81] G. W. Stewart. Matrix calculations on hand-held calculators. ACM SIGNUM Newsletter, 16(1):10–13, March 1981. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Stewart:1991:FFT

[Ste91] G. W. Stewart. FTP — file transfer program. ACM SIGNUM Newsletter, 26(4):2–3, October 1991. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Stetter:1996:NCA

[Ste96a] Hans J. Stetter. Numerical computer algebra. ACM SIGNUM Newsletter, 31(3):2–??, July 1996. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Stetter:1996:SNA

[Ste96b] Hans J. Stetter. Symbolic numeric algebra for polynomials. ACM SIGNUM Newsletter, 31(4):2–??, October 1996. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Strehmel:1976:NCM

[Str76] Karl Strehmel. A new class of multistep methods for systems of first order differential equations with exponential adaptation. ACM SIGNUM Newsletter, 11(1):46–47, May 1976. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Strehmel:1977:NAN

[Str77] Karl Strehmel. A new algorithm for the numerical solution of ordinary stiff differential equations of the nth order. ACM SIGNUM Newsletter, 12(4):20–22, December 1977. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Streit:1981:SLH

[Str81] Roy L. Streit. Solution of large Hermitian eigenproblems on virtual and cache memory computers. ACM SIGNUM Newsletter, 16(2):6– 7, June 1981. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Sundstrom:1975:EIP

[SW75] Arne Sundström and Lars Westerling. ENEP: an interactive program for solving one-dimensional continuum mechanical problems using a graphic terminal (abstract). ACM SIGNUM Newsletter, 10(4):38–39, December 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Tang:1975:ASL

[Tan75] I. Tang. An algorithm for a system of linear equations. ACM SIGNUM Newsletter, 10(2–3):42–44, November 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Tweedale:1975:DUI

[TC75] A. Tweedale and C. Chan. On the design of a user interface to large FORTRAN subroutine packages under IBM 360/370 OS. ACM SIGNUM Newsletter, 10(1):15–17, January 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic). See note [Boy75].

Tee:1972:SEI

[Tee72] Garry J. Tee. A simple example of an ill-conditioned matrix. ACM SIGNUM Newsletter, 7(3): 19–20, October 1972. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Tee:1974:EZM

[Tee74] Garry J. Tee. The eigenvectors of Zielke’s matrices. ACM SIGNUM Newsletter, 9(2):23, April 1974. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Tee:1980:KEA

[Tee80] G. J. Tee. Khachian’s efficient algorithm for linear inequalities and linear programming. ACM SIGNUM Newsletter, 15(1):13–15, March 1980. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Temperton:1983:CVC

[Tem83] Clive Temperton. Cray 1 v. Cyber 205: some comparisons. ACM SIGNUM Newsletter, 18(2):20–21, April 1983. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Thompson:1987:IEF

[Tho87] Peter Thompson. Implementing an elementary function library. ACM SIGNUM Newsletter, 22(2):2– 5, April 1987. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

TrejoRamirez:1997:HRN

[TK97] Raúl Antonio Trejo Ramírez and Vladik Kreinovich. How to rate numerical packages? Foundations of ratings, and their possible use in numerical mathematics, in education, and in military design. ACM SIGNUM Newsletter, 32(1–2): 2–10, January/April 1997. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Traub:1966:PDQ

[Tra66] J. F. Traub. Ph.D. qualifying examination: numerical analysis. ACM SIGNUM Newsletter, 1(2):??, July 1966. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Traub:1967:EQQ

[Tra67a] J. F. Traub. Examination questions: qualifying examination: numerical analysis. ACM SIGNUM Newsletter, 1(3):9–11, February 1967. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Traub:1967:NAA

[Tra67b] J. F. Traub. Numerical analysis: approximation or estimation? ACM SIGNUM Newsletter, 2(2):??, June 1967. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Traub:1967:PDQ

[Tra67c] J. F. Traub. Ph.D. qualifying examination: numerical analysis. ACM SIGNUM Newsletter, 2(3):??, December 1967. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Traub:1968:SAC

[Tra68] J. F. Traub. State of the art in calculation of polynomial zeros. ACM SIGNUM Newsletter, 3(3):??, July 1968. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Trefethen:1985:TMG

[Tre85] Lloyd N. Trefethen. Three mysteries of Gaussian elimination. ACM SIGNUM Newsletter, 20(4):2–5, October 1985. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Turner:1969:DSC

[Tur69] L. Richard Turner. Difficulty in sin/cos routine. ACM SIGNUM Newsletter, 4(3):13, October 1969. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Tydeman:1994:WCT

[Tyd94] Fred Tydeman. What causes a trap in IEEE-754 floating-point? ACM SIGNUM Newsletter, 29(1): 2–4, January 1994. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Usow:1969:GEA

[UF69] Karl H. Usow and Lloyd D. Fosdick. Guidelines for evaluating an algorithm for publication. ACM SIGNUM Newsletter, 4(3):19–20, October 1969. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Usow:1969:SSC

[Uso69] Karl H. Usow. SIGNUM subroutine certification committee. ACM SIGNUM Newsletter, 4(3):15–18, October 1969. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Usow:1970:CB

[Uso70] K. H. Usow. Certification bibliography. ACM SIGNUM Newsletter, 5(2):14–15, March 1970. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Vandevender:1982:SMS

[VH82] W. H. Vandevender and K. H. Haskell. The SLATEC mathematical subroutine library. ACM SIGNUM Newsletter, 17(3):16– 21, September 1982. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Voigt:1981:SRA

[Voi81] Robert G. Voigt. Summary of research activity at ICASE. ACM SIGNUM Newsletter, 16(3):12– 13, September 1981. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Volper:1979:UPN

[Vol79] Dennis Volper. Using Pascal for numerical analysis. In ACM [ACM79], pages 93–95.

Wagener:1984:SWT

[Wag84] Jerrold L. Wagener. Status of work toward revision of programming language Fortran. ACM SIGNUM Newsletter, 19(3):1–42, July 1984.

CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Wang:1978:EPF

[Wan78] J. Y. Wang. The evaluation of periodic functions with large input arguments. ACM SIGNUM Newsletter, 13(4):7–8, December 1978. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Wanner:1979:SOM

[Wan79] Gerhard Wanner. A-stability of oneand multistep methods. In Skeel [Ske79b], page ??

Warner:1976:NVO

[War76] D. D. Warner. A note on variable order strategies for differential equation solvers. ACM SIGNUM Newsletter, 11(3):27–28, October 1976. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Warming:1979:FSL

[WB79] R. F. Warming and Richard M.

Beam. Factored, A-stable, linear multistep methods — an alternative to the method of lines for multidimensions. ACM SIGNUM Newsletter, 14(3):17–19, September 1979. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Ward:1980:NLA

[WG80] Robert C. Ward and Gene H.

Golub. Numerical linear algebra. ACM SIGNUM Newsletter, 15(3): 9–26, September 1980. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Whiten:1994:DPR

[Whi94] W. J. Whiten. Determination of parameter relations within non-linear models. ACM SIGNUM Newsletter, 29(3–4):2–5, July/October 1994. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Williams:1986:MSE

[Wil86] Donald L. Williams. A mathematical software environment. ACM SIGNUM Newsletter, 21(3):2–12, July 1986. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Wisniewski:1978:SEC

[Wis78] John A. Wisniewski. Some experiments with computing the complex absolute value. ACM SIGNUM Newsletter, 13(1):11–12, March 1978. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Witte:1967:CTT

[Wit67] Bruno F. W. Witte. Comparative tabulations of test problems and test results for mathematical computer routines. ACM SIGNUM Newsletter, 2(3):??, December 1967. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Whiten:1995:RTF

[WK95] W. J. Whiten and L. Kocis. Regarding test functions for multidimensional integration. ACM SIGNUM Newsletter, 30(1):2–7, January 1995. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Walster:1996:UBE

[WK96] G. William Walster and Vladik Kreinovich. For unknown-but-bounded errors, interval estimates are often better than averaging. ACM SIGNUM Newsletter, 31(2):6– 19, April 1996. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Xanthis:1986:POM

[Xan86] Leonidas S. Xanthis. A pseudo-ODE modelling trick for the direct method of lines computation of important fracture mechanics parameters. ACM SIGNUM Newsletter, 21(1–2):10–16, April 1986. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Zahir:1989:CAD

[Zah89] M. S. Zahir. A computational algorithm for LU decomposition of a sequence of matrices that differ only in one or more elements. ACM SIGNUM Newsletter, 24(4):8–16, October 1989. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Zielke:1973:NTM

[Zie73] Gerhard Zielke. A new test matrix for inverting matrices. ACM SIGNUM Newsletter, 8(2):22–23, April 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Zielke:1974:GTM

[Zie74] Gerhard Zielke. Generalization of Tee’s matrix. ACM SIGNUM Newsletter, 9(1):11–12, January 1974. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Zielke:1978:IRG

[Zie78a] Gerhard Zielke. Iterative refinement of generalized matrix inverses now practicable. ACM SIGNUM Newsletter, 13(4):9–10, December 1978. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Zielke:1978:TMG

[Zie78b] Gerhard Zielke. Test matrices for generalized inverses. ACM SIGNUM Newsletter, 13(4):10–12, December 1978. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Zielke:1981:GRT

[Zie81] Gerhard Zielke. Generalizations of a Rutishauser test matrix with exact Moore–Penrose inverses. ACM SIGNUM Newsletter, 16(3): 7–8, September 1981. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).